### v1

# Robust Radiotherapy Planning with Spatially Based Uncertainty Sets

Abstract
Radiotherapy treatment planning is a challenging large-scale optimization problem plagued by uncertainty. Following the robust optimization methodology, we propose a novel, spatially based uncertainty set for robust modeling of radiotherapy planning, producing solutions that are immune to unexpected changes in biological conditions. Our proposed uncertainty set realistically captures biological radiosensitivity patterns that are observed using recent advances in imaging, while its parameters can be personalized for individual patients. We exploit the structure of this set to devise a compact reformulation of the robust model. We develop a row-generation scheme to solve real, large-scale instances of the robust model. This method is then extended to a relaxation-based scheme for enforcing challenging, yet clinically important, dose-volume cardinality constraints. The computational performance of our algorithms, as well as the quality and robustness of the computed treatment plans, are demonstrated on simulated and real imaging data. Based on accepted performance measures, such as minimal target dose and homogeneity, these examples demonstrate that the spatially robust model achieves almost the same performance as the nominal model in the nominal scenario, and otherwise, the spatial model outperforms both the nominal and the box-uncertainty models.
Keywords: Radiotherapy planning, robust optimization, biomarker uncertainty, row and column generation
1 Introduction
Radiotherapy treatment, and in particular intensity modulated radiation therapy (IMRT), involves setting up an array of beams with variable intensities to irradiate a tumor. Planning such a treatment entails selecting the intensity of each beam or beam unit (called a beamlet). To facilitate such planning, the effect of the beam intensities – that is, the resulting radiation dose (whose unit of measure is 1 Gy = 1 J/kg) – on both the tumor and the healthy surrounding organs must be analyzed. The dose is measured over a three-dimensional grid of spatial cube units called voxels. The size of a voxel, and the granularity of the partition into voxels, are determined by the radiation beamlet size and the imaging resolution. Recent technology allows for voxel edge lengths of 2-3 mm. The dose at each voxel is generally assumed to be linearly related to the chosen beamlet intensities. In general, this is the theme of radiotherapy treatment planning (RTP), which may prescribe the treatment over both time and space. The particular optimization problem for setting up the beamlet intensities is known as fluence map optimization (FMO). A variety of formulations have been proposed for this problem. Typically, such formulations require that the tumor voxels are sufficiently irradiated, while healthy tissue voxels are spared from receiving an unhealthy dose. The formulations differ in the way that these requirements are handled. Additional clinical requirements can also be factored into the model. Some formulations aim to maximize the average tumor voxel dose (Preciado-Walters et al. 2006), while constraining the dose received by healthy organs. Other formulations minimize a measure of the dose received by healthy tissue voxels, or by all voxels, subject to ensuring that target (i.e., tumor) voxels receive at least the prescribed dose (Romeijn et al. 2006).
Our approach is to maximize the tumor’s “weakest link”: the voxels receiving the minimum dose of all tumor voxels. This is a widely accepted objective for evaluating radiotherapy planning (Deasy 1997, Yan et al. 1997, ICR 2010, Preciado-Walters et al. 2004), subject to constraints on the healthy organ dose and on tumor dose homogeneity. In particular, the objective of maximizing the minimum dose objective is argued to better support biological optimization, as “a conservative surrogate for biological effect”, compared with minimization of deviations from prescribed doses (Deasy 1997). The importance of combining this objective with homogeneity constraints is outlined by Yan et al. (1997), Allen Li et al. (2012), Khairi et al. (2021).
Advances in imaging have exposed uncertainties in the effect of the prescribed radiation dosage. It is now known that tumors are not homogeneous, but rather, the sensitivity of cells varies across the tumor and changes over time (Bentzen and Gregoire 2011, Saka et al. 2014, Titz and Jeraj 2008, Nohadani and Roy 2017). In an effort to improve patient outcomes, there is a need for the model to account for the uncertain problem data. Modeling uncertainty in an optimization problem usually results in added complexity and scale. Uncertainty is involved in all radiotherapy treatment planning stages, starting with the target volume definition, but also in relation to the dose required to control the tumor, healthy tissue tolerances, and the actual dose delivered to the patient. Uncertainty in radiotherapy planning has been extensively addressed in the literature, mainly in the context of geometric uncertainty that results from the patient’s movement and from imaging inaccuracies; see, for example, Bortfeld et al. (2008), the recent survey by Unkelbach et al. (2018), and the references therein.
Approaches for dealing with uncertainty in FMO include stochastic programming, robust optimization and distributionally robust optimization. The stochastic approach assumes that the distribution of the uncertainty is known and its parameters can be reasonably estimated. In practice, estimating the exact distribution is difficult, and incorrectly estimating the distribution may lead to solutions that are not clinically viable. In contrast, the robust optimization methodology (Ben-Tal et al. 2009) does not assume any knowledge of the distribution, but rather, defines a set that contains all possible uncertainty realizations that need to be hedged against. By making few or no assumptions about the distribution, the robust approach optimizes the treatment for the worst-case realization of the uncertainty. Motivated by motion uncertainty, Chu et al. (2006) propose a robust formulation with uncertain influence matrices. The distributionally robust approach combines the stochastic and robust approaches – it protects against the worst-case distribution within a family of possible probability distributions of the uncertain parameters. For example, Bortfeld et al. (2008) model motion uncertainty by defining families of distributions that correspond to possible motion patterns of patients. Unkelbach et al. (2007) consider a (discrete) uncertainty set for beamlet range uncertainties in proton therapy (IMPT). This uncertainty set is constructed so that beamlets of the same beam must reach the same range. In this work they consider the objective of minimizing a weighted sum of the deviations from a prescribed dose. However, the model being solved may be overly conservative as the worst case is applied to each of the voxels separately, rather than to the objective. A similar scenario-based approach to account for range uncertainty and setup uncertainty in IMPT is also used in Pflugfelder et al. (2008).
Another source of uncertainty, also mentioned in Unkelbach et al. (2018), arises from the biological conditions of the patients, which the authors consider a new frontier of robust RTP optimization. These conditions may affect the radiosensitivity of the tumor to the dose. The resulting effective dose is referred to as the biologically-adjusted dose. We note that this adjustment is different from the biological effective dose (BED) often modeled by the linear-quadratic (LQ) model (Jones et al. 2001), which aims to estimate the biological effectiveness in terms of cell-kill probability for a given physical dose. In contrast, the biologically adjusted dose corrects the physical dose to account for the impact of a biological condition on the effective dose at a particular voxel. For the case of oxygen levels (or the lack thereof also known as hypoxia), this is described by the oxygen enhancement ratio (OER), which is the ratio between doses required to obtain the same biological effect (Hill et al. 2015, McMahon 2018), or its normalized variant known as the oxygen modification factor (OMF) (Titz and Jeraj 2008). Information about a living tissue’s (e.g., a tumor’s) biological conditions is available through bio-markers and imaging data. Specifically, the tumor’s oxygenation level, which strongly affects its radiosensitivity, is measured by imaging techniques such as fluoromisonidazole (FMISO) (Toma-Dasu et al. 2012). Translating the imaging data into radiosensitivity values is a crude estimation that is subject to uncertainty. Further, the oxygenation level, and accordingly the radiosensitivity, may change between the time of the image scan and the commencement of treatment. In fact, biological uncertainty in radiation therapy, and the effect of oxygenation on radiosensitivity in particular, is an active area of research in the medical physics and other research communities (Titz and Jeraj 2008, Toma-Dasu et al. 2012, Shang et al. 2021). Biological conditions and radiosensitivities are considered in the context of FMO by Saka et al. (2014) but their proposed methods do not attempt to model and protect from the associated uncertainty. Robust static (e.g., cumulative) dose planning under biological uncertainty has been proposed for biological uncertainty by Li et al. (2015) and Ten Eikelder et al. (2020). Ten Eikelder et al. (2020) propose a tumor control probability (TCP) based model assuming voxel independence and simple box uncertainty for its parameters. Simple box uncertainty here implies conservative solutions where all voxels are assumed to be simultaneously insensitive. Li et al. (2015) propose a dose-based model with budgeted uncertainty (see Bertsimas and Sim (2004)) on the deviation of the voxels’ prescribed doses from their nominal value. In this model, the conservatism is controlled by a single budget parameter that indicates how many voxel doses can deviate. However, in this model the budget is allocated linearly to the change in the prescribed doses, so it might not accurately capture measurement errors whose effect on radiosensitivity may be highly nonlinear Titz and Jeraj (2008). Also, this model does not limit the differences in radiosensitivities among neighboring voxels.
Recent papers that consider radiosensitivity uncertainty in adaptive radiation therapy planning include Nohadani and Roy (2017), Roy et al. (2022), Jeyakumar et al. (2023). These papers assume a perscribed dose approach where the objective is to minimize the total dose administered while satisfying simple lower and upper dose bound constraints on target voxels and OARs, respectively. Consequently, the biological uncertainty in this formulation appears only on the right-hand side. Nohadani and Roy (2017) utilize simple box uncertainty sets containing the measured radiosensitivities, and model the evolution of this uncertainty through time. In the current paper, we also formulate and solve a robust FMO that accounts for biological uncertainty at the voxel level where the main focus is to model the spatial connectivity or dependence of the voxel uncertainties. Both Roy et al. (2022), Jeyakumar et al. (2023) propose models similar to Nohadani and Roy (2017). Roy et al. (2022) studies an extension of the model to a multistage adaptive problem, while Jeyakumar et al. (2023) consider adaptivity to uncertainty in influence matrix resulting in a semidefinite program.
In contrast to previous work, our approach is to model and explicitly account for the spatial interdependence among individual voxel radiosenstivity uncertainties. Accounting for such interdependence is expected to improve over simpler robust approaches that tend to result in overly conservative solutions. Spatial dependence has been considered in other applications; for example by estimating covariance matrices and applying these using ellipsoid uncertainty sets in transportation Chassein et al. (2019), or general polyhedral uncertainty sets in energy Lorca and Sun (2014). These previous approaches may be less appropriate for biological uncertainty in radiotherapy because there may not be sufficient data to reliability estimate the correlations between voxel radiosensitivities. Also, applying both types of uncertainty sets to large-scale radiotherapy optimization problems, may be computationally intractable in practice. In radiotherapy, the scenario-based approaches to range and setup uncertainties in IMPT appear to resemble robust optimization with spatial uncertainty (for example Unkelbach et al. (2007), Pflugfelder et al. (2008) discussed above). However, the uncertainty applies to the beamlets rather than to voxels and it is addressed by scenario-based approaches, so the dependency is not explicitly modeled. In contrast, the current paper focuses on spatial uncertainty modelling, reformulation of the corresponding robust FMO problem and solution methods for solving this problem. Dividing the administered dose into fractions and accounting for treatment sessions over time (also known as fractions) may involve a more elaborate uncertainty model over both space and time that is left for future work.
To make our proposed model applicable to clinical settings, we also consider important dose-volume constraints. These constraints limit the percentage of healthy-organ voxels that can receive more than a specified radiation dose. These are hard combinatorial constraints, which are virtually impossible to handle exactly given the problem size. Hence, many solution approaches have been proposed to approximately solve such constraints in the literature (Ferris et al. 2003, Preciado-Walters et al. 2006, Romeijn et al. 2006). However, effectively solving models that incorporate such constraints remains a challenging problem, especially in the context of the large scale problems considered herein.
To summarize our current contribution, we propose a novel FMO formulation that accounts for voxel radiosensitivity and its associated uncertainty, in both the objective and in the homogeneity constraints. To this end, Section 2 presents our novel spatially dependent uncertainty set with justification based on publicly available real biomarker (FMISO imaging) data. In Section 3, we develop a scalable constraint generation algorithm for solving the resulting robust formulation. Section 4 presents a method for relaxing the dose-volume constraints – a method that can be solved and tuned in order to determine a feasible solution to the robust formulation when augmented with a dose-volume constraint. Finally, Section 5 includes a numerical study that demonstrates the scalability of our method, shows how to personalize the treatment plan, and presents a study of the effectiveness of our method under different scenarios.
2 Modeling Robust Radiotherapy Planning
We now present our model for robust radiotherapy planning with biological uncertainty. In Section 2.1, the nominal model, which is the planning optimization model in the absence of uncertainty, is discussed. In Section 2.2, the robust approach to hedging against uncertainty is presented. This approach includes the development of a novel biological uncertainty set and its incorporation into the nominal problem.
2.1 Model Introduction - Nominal Model
In the following, we use the notation $[n]=\{1,\ldots,n\}$, for any integer $n\geq 1$. Additionally, for a vector $a\in\mathbb{R}^{n}$ and index set $I\subset[n]$, $a_{I}\in\mathbb{R}^{\left|I\right|}$ denotes the subvector of $a$ with coordinate set $I$. The set of all voxel indices is $[m]$ for some positive integer $m$, which is partitioned into the planning target volume (PTV) index set $T\subset[m]$, containing tumor voxel indices, and index sets corresponding to the $K$ organs at risk (OARs), $H_{k}\subset[m]$, for each $k\in[K]$, such that $[m]=T\cup\bigcup_{k=1}^{K}H_{k}$.
FMO involves determination of the $n$ beamlet intensities. For $i\in[n]$, $x_{i}$ is beamlet $i$’s intensity decision variable. Given the intensity vector ${\bf x}\in\mathbb{R}^{n}$ and an influence matrix ${\bf D}\in\mathbb{R}^{m\times n}$, for each voxel $v\in[m]$, $d_{v}(x)=\sum_{i=1}^{n}D_{vi}x_{i}$ is the physical dose applied to voxel $v$.
The objective is to maximize a concave function of the dose associated with the PTV voxels, $f({\bf d}_{T})$, while constraining the OAR dose. For convenience, throughout most of the paper, we will consider a simple OAR bound of the form $d_{v}(x)\leq\bar{d}_{k}$ for each $k\in[K]$ and $v\in H_{k}$, where $\bar{d}_{k}$ is a predefined upper bound on the dose for OAR $k\in[K]$. In Section 4, we consider more elaborate dose-volume constraints. An important aspect of the problem that affects the definition of the function $f$, which is also the main focus of the model considered in the current paper, is that the effect of the dose on a given PTV voxel $v\in T$ depends on biological factors, such as the PTV’s oxygenation level. This information is modeled by a radiosensitivity vector $\phi\in[0,1]^{\left|T\right|}$. Accordingly, the biologically adjusted dose (or simply adjusted dose) of voxel $v$ is given by $\phi_{v}{d}_{v}$. For example, if considering the effect of oxygen levels then $\phi$ would be a vector of OMF values, as discussed in Section 1. The radiosensitivity is usually computed based on biomarker and imaging data. Notably, $\phi$ is highly uncertain due to imaging and conversion inaccuracies and variation in the oxygen level over time. Yet, the radiotherapy planning goal is generally to administer as much adjusted dose to the PTV voxels as possible, which is accomplished here by setting $f({\bf d})=\min_{v\in T}\phi_{v}d_{v}(x)$.
We also consider homogeneity constraints that limit the dose range across the PTV (Preciado-Walters et al. 2004). This is intended to prevent the creation of “hot spots” in the tumor tissue, which cause damage to the healthy tissue in which the tumor is embedded. Specifically, our homogeneity constraints are of the form $\mu f({\bf d})-\phi_{v}{d}_{v}\geq 0$, for all $v\in T$, where $\mu>1$ is a given homogeneity parameter. Homogeneity has been traditionally applied to the physical dose Allen Li et al. (2012), Kataria et al. (2012), Khairi et al. (2021). The possibility of applying it to the adjusted dose has been enabled by recent technological advances in estimating necessary biological information, specifically the radiosensitivity vector $\phi$. The main goal of the current paper is to address the challenge of modeling the uncertainty in $\phi$ and using this information to improve radiotherapy treatment planning. The optimization problem of determining the intensity of the beams given the radiosensitivity parameter $\boldsymbol{\phi}$ is formulated as
$\underset{\begin{subarray}{c}{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R}\end{subarray}}{\text{maximum}}$ $\underaccent{\bar}{d}$ (1b)
subject to $\phi_{v}d_{v}(x)\geq\underaccent{\bar}{d},$ $v\in T$ (1c)
$\mu\underaccent{\bar}{d}-\phi_{v}d_{v}(x)\geq 0,$ $v\in T$ (1d)
$d_{v}(x)\leq\bar{d}_{k},$ $k\in[K],v\in H_{k},$ (1e)
which will be referred to as the nominal formulation.
Note that $\boldsymbol{\phi}$ is derived from imaging data through a series of crude transformations. Thus, it is potentially highly inaccurate. Ignoring this inaccuracy in practice may lead to poor clinical outcomes in terms of controlling the adjusted dose and adjusted homogeneity. The next section extends the model to address the uncertainty in $\boldsymbol{\phi}$.
2.2 Robust Optimization Model with Spatially Bound Uncertainty Set
We now present a modeling scheme to deal with the uncertainty in the parameter vector $\boldsymbol{\phi}$. The radiosensitivity uncertainty is caused by multiple factors, including imaging noise, biomarker conversion error, and variation over time. Inferring a distribution of the radiosensitivity parameter based on limited patient-specific (personalized) data is nontrivial and prone to error. Moreover, maximizing the clinical outcome (in our model, maximizing the minimum PTV adjusted dose) while using a high-dimensional multivariate probability distribution to model the biological uncertainty, would result in chance-constrained optimization problem (Luedtke et al. 2010), which is notorious for being computationally hard. Instead, our approach to modeling uncertainty follows the robust optimization (RO) methodology. In this modeling framework, the uncertain parameters are assumed to lie in a predetermined uncertainty set denoted by $\mathcal{U}$. An RO problem formulation is designed to protect its solutions against all realizations in the uncertainty set. In particular, an optimal solution $({\bf x},\underaccent{\bar}{d})$ must remain feasible for every possible realization $\boldsymbol{\phi}\in\mathcal{U}$ while maximizing the worst-case (minimum) PTV adjusted dose.
Formulating the robust counterpart (RC) of (2.1), in which variable $\underline{d}$ now depends on the uncertain $\phi$ subject to uncertain homogeneity constraints (1d), would lead to an adjustable RO problem. Such problems are computationally intensive (Ben-Tal et al. 2004), and although it may be possible to solve these adjustable formulations when the problem is moderately sized, we find it more effective to avoid the adjustable robust model by reformulating the homogeneity constraints according to the equivalence
$\mu\phi_{u}{d}_{u}-\phi_{v}{d}_{v}\geq 0,\text{ for all }v,u\in T$$\Leftrightarrow$ $\mu f({\bf d})-\phi_{v}{d}_{v}\geq 0\text{, for all }v\in T,$ (2)
at the cost of increasing the number of constraints by a factor of $O(\left|T\right|)$. The resulting robust problem is given by
$\underset{\begin{subarray}{c}{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R}\end{subarray}}{\text{maximum}}$ $\underaccent{\bar}{d}$ (3b)
subject to $\min_{\boldsymbol{\phi}\in\mathcal{U}}\phi_{v}d_{v}(x)\geq\underaccent{\bar}{d},$ $v\in T$ (3c)
$\max_{\boldsymbol{\phi}\in\mathcal{U}}\left\{\phi_{v}d_{v}(x)-\mu\phi_{u}d_{u}(x)\right\}\leq 0,$ $u,v\in T$ (3d)
$d_{v}(x)\leq\bar{d}_{k},$ $k\in[K],v\in H_{k}.$ (3e)
Constructing the uncertainty set poses some challenges, in particular, trying to satisfy and balance the following set of requirements for $\mathcal{U}$:
  1. (i)
    Personalized for the patient. This requirement relies on patient specific information, such as imaging data acquired at the start of the treatment.
  2. (ii)
    Computationally tractable. The nominal problem may already be computationally intensive given the large number of voxels. The computational effort of solving the robust problem should not be excessive; in particular, it should not significantly increase the running time relative to that required to solve the nominal problem.
  3. (iii)
    Not too conservative. A goal of our modeling approach is to determine a set that captures the essence of the uncertainty, thereby providing “reasonable” protection, while not being too conservative, and thus tackling one of the major criticisms of RO.
In the next section, we present a novel uncertainty set that models the spatial dependence between voxels while addressing these requirements.
2.2.1 The Spatially Bound Uncertainty Set.
Given a radiosensitivity vector $\widehat{\boldsymbol{\phi}}$ that is estimated based on a patient’s imaging data at the start of the treatment, the goal is to construct an uncertainty set that satisfies our three modeling requirements, described in Section 2.2. Perhaps the simplest uncertainty set that can be applied is a box (in particular, a hypercube) uncertainty set, which allows the radiosensitivity of each PTV voxel $v$ to deviate from its calculated value $\widehat{\phi}_{v}$ by at most $\delta>0$, where $\delta$ is a given parameter. That is, $\mathcal{U}_{\text{B}}=\{\boldsymbol{\phi}\in\mathbb{R}^{|T|}:|\phi_{v}-\widehat{\phi}_{v}|\leq\delta,\;\forall v\in T\}$. This uncertainty set, used for example in Nohadani and Roy (2017), admits a straightforward robust formulation. However, if $\delta$ is chosen to be sufficiently large to capture the actual uncertainty in the tumor biological conditions, then the resulting set may be overly conservative, thereby violating one of our key modeling requirements (see Section 2.2). To restrict the uncertainty further, a key observation to be made is that voxels that are situated close to each other should have similar radiosensitivities. In Figure 1, both plots demonstrate that the radiosensitivity differences are small for neighboring voxels and increase with the distance.
A Pairwise radiosensitivity difference percentiles vs. pairwise distances.
B Histogram of the pairwise radiosensitivity differences for pairwise distances of 1, 2, 3, and all.
Figure 1: Relationship between radiosensitivity difference and pairwise PTV voxel distance in the first-visit data of Patient 1 from the TCIA brain dataset (Clark et al. 2013).
Incorporating this observation into the uncertainty model leads to a spatially bound (SB) uncertainty set; in particular, we define
$\mathcal{U}_{\text{SB}}=\{\boldsymbol{\phi}\in\mathbb{R}^{|T|}:|\phi_{v}-\widehat{\phi}_{v}|\leq\delta,\;|\phi_{v}-\phi_{u}|\leq\gamma_{vu},\;\forall v,u\in T\},$ (4)
where $\gamma_{vu}$ are parameters of the uncertainty set that may depend on characteristics of the voxel pair $v$ and $u$. Specifically, the $\gamma_{uv}$ parameters are assumed to depend on the pairwise distances, and in particular, they are given by a metric-preserving function of the distance.
Assumption 1.
For any $u,v\in T$, $\gamma_{uv}=\Gamma(\operatorname{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dist}}(u,v))$, where $\operatorname{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dist}}(u,v)$ is the distance in voxels between $u$ and $v$, and $\Gamma:\mathbb{R}\rightarrow[0,1]$ is a nondecreasing subadditive function satisfying $\Gamma(x)=0$ if and only if $x=0$.
Defining $\gamma_{uv}$ as a function of the distance between $u$ and $v$ in this way captures the interdependence of the radiosensitivities of nearby voxels and limits the conservatism associated with the box uncertainty set. To the best of our knowledge, such spatially dependent uncertainty sets have not been previously investigated in the context of robust radiotherapy planning, nor it seems in the context of RO in general.
The $\gamma_{uv}$ parameters for $u,v\in T$ can be personalized to a particular patient’s data, thereby addressing requirement (i). Requirement (iii) may be satisfied by imposing the spatial constraints associated with the parameters $\gamma_{uv}$, for $u,v\in T$, to augment the box constraints in the definition of the spatially bound uncertainty set. In the absence of these spatial constraints, the uncertainty set would reduce to the overly conservative box uncertainty set. The following section addresses the question of whether optimal solutions of (2.2) for uncertainty set $\mathcal{U}_{\text{SB}}$ can be computed efficiently in practice (requirement (ii)).
2.2.2 Compactly Reformulating The Robust Problem
In this section, we present an exact reformulation of (2.2) for the case where $\mathcal{U}=\mathcal{U}_{\text{SB}}$ with a polynomial number of constraints with respect to the size of the nominal formulation (2.1).
The following result, which will be used in the proofs that follow, establishes that $\gamma_{uv}$ is a metric.
Lemma 2 (Corazza 1999, Proposition 3.2).
Suppose that Assumption 1 holds. Then, $\gamma_{\cdot}=\Gamma(d(\cdot)):T\times T\rightarrow[0,1]$ is a metric.
For $S\subseteq\mathbb{R}^{n}$, denote the projection of set $S$ onto index set $I\subseteq[n]$ by $P_{I}(S)$, so that $P_{I}(S)=\left\{v_{I}\;\left|\;\;v\in S\right.\right\}$. For simplicity, for the singleton $I=\{i\}$, the projection will be denoted by $P_{i}(S)$. The following is a straightforward observation that is used in reformulating (2.2).
Observation 3.
Let $\mathcal{U}\subseteq\mathbb{R}^{n}$ be a compact set. Then, for any continuous function $f:\mathbb{R}^{|I|}\rightarrow\mathbb{R}$, $\min_{{\bf u}\in\mathcal{U}}f({\bf u}_{I})=\min_{{\bf v}\in P_{I}(\mathcal{U})}f({\bf v})$.
Problem (2.2) is now specialized and reformulated for the uncertainty set $\mathcal{U}_{\text{SB}}$. Henceforth, it is assumed that $\mathcal{U}_{\text{SB}}\neq\emptyset$. Starting with the the first set of constraints (3c), and using Observation 3, for each $v\in T$, the inner minimization in this constraint is given by
$\min_{\boldsymbol{\phi}\in\mathcal{U}_{\text{SB}}}\phi_{v}d_{v}(x)=\min_{\phi_{v}\in P_{v}(\mathcal{U}_{\text{SB}})}\phi_{v}d_{v}(x).$ (5)
Since matrix ${\bf D}$ and vector ${\bf x}$ are both nonnegative, the solution of (5) is given by the minimal value of $\phi_{v}$. For convenience, in the following, for each $u\in T$, define $\underaccent{\bar}{\phi}^{0}_{u}=\max\{0,\widehat{\phi}_{u}-\delta\}$, $\bar{\phi}^{0}_{u}=\min\{1,\widehat{\phi}_{u}+\delta\}$, $\underaccent{\bar}{\phi}_{u}=\min_{\boldsymbol{\phi}\in\mathcal{U}_{\text{SB}}}\phi_{u}$, and $\bar{\phi}_{u}=\max_{\boldsymbol{\phi}\in\mathcal{U}_{\text{SB}}}\phi_{u}$. The following proposition characterizes the one-dimensional projection of $\mathcal{U}_{\text{SB}}$.
$\phi$$\hat{\phi}_{v}$|$v$|$u$$\hat{\phi}_{u}$|$w$$\hat{\phi}_{w}$$\bar{\phi}^{0}_{v}$$\underaccent{\bar}{\phi}^{0}_{v}$$\bar{\phi}^{0}_{u}$$\underaccent{\bar}{\phi}^{0}_{u}$$\bar{\phi}^{0}_{w}$$\underaccent{\bar}{\phi}^{0}_{w}$$+\bar{\gamma}_{uv}$$-\bar{\gamma}_{uv}$$+\bar{\gamma}_{wv}$$-\bar{\gamma}_{wv}$$\bar{\phi}_{v}$$\underaccent{\bar}{\phi}_{v}$
Figure 2: An illustration of the one-dimensional projection result of Proposition 4. Assuming only three voxels, $v,u,$ and $w$, the bounds on the radiosenstivity of voxel $v$ are determined by tightening its original interval ${[{\underaccent{\bar}{\phi}}^{0}_{v},\bar{\phi}^{0}_{v}]}$ by the intersection with $[\underaccent{\bar}{\phi}^{0}_{u}-\gamma_{uv},\bar{\phi}^{0}_{u}+\gamma_{uv}]$ and $[\underaccent{\bar}{\phi}^{0}_{w}-\gamma_{wv},\bar{\phi}^{0}_{w}+\gamma_{wv}]$, imposed by voxels $u$ and $w$, respectively.
Proposition 4.
For every $v\in T$:
  1. (i)
[…]
  2. (ii)
    $\underaccent{\bar}{\phi}_{v}\geq\max_{u\in T\setminus\{v\}}\{\underaccent{\bar}{\phi}_{u}-\gamma_{uv}\}$ and $\bar{\phi}_{v}\leq\min_{u\in T\setminus\{v\}}\{\bar{\phi}_{u}+\gamma_{uv}\}$.
  3. (iii)
    Suppose that Assumption 1 holds. Then,
    $\underaccent{\bar}{\phi}_{v}=\max_{u\in T}\{\underaccent{\bar}{\phi}^{0}_{u}-\gamma_{uv}\}$ and $\bar{\phi}_{v}=\min_{u\in T}\{\bar{\phi}^{0}_{u}+\gamma_{vu}\}.$ (6)
Proof.
Proof.Part (i): $P_{v}(\mathcal{U}_{\text{SB}})=[\underaccent{\bar}{\phi}_{v},\bar{\phi}_{v}]$ follows from that fact that $\mathcal{U}_{\text{SB}}$ is a convex and compact set and from Observation 3, which together imply that the projection of $\mathcal{U}_{\text{SB}}$ onto $v$ is a closed interval $[\underaccent{\bar}{\phi}_{v},\bar{\phi}_{v}]$.
Part (ii): For the sake of contradiction, suppose that $\underaccent{\bar}{\phi}_{v}<\underaccent{\bar}{\phi}_{u}-\gamma_{uv}$ for some $u\in T\setminus\{v\}$, and let $\phi^{*}$ be a minimizer (vector) that attains $\underaccent{\bar}{\phi}_{v}$. Then, by optimality of $\phi^{*}_{v}=\underaccent{\bar}{\phi}_{v}$, $\phi^{*}_{v}-\phi^{*}_{u}\leq\underaccent{\bar}{\phi}_{v}-\underaccent{\bar}{\phi}_{u}<-\gamma_{uv}$, thereby contradicting the feasibility of $\phi^{*}$. The proof of the upper bound on $\bar{\phi}_{v}$ is similar.
Part (iii): Let $\phi^{*}\in\operatorname{argmax}_{\phi\in\mathcal{U}_{\text{SB}}}\{\phi_{v}\}$, so that $\bar{\phi}_{v}=\phi^{*}_{v}$. First, observe that by the optimality of $\phi^{*}_{v}$, there is at least one binding constraint, $\bar{\phi}_{v}\leq\bar{\phi}^{0}_{v}$ or $\bar{\phi}_{v}-\phi^{*}_{u_{1}}\leq\gamma_{vu_{1}}$, for some $u_{1}\in T$. Thus, $\bar{\phi}_{v}=\min\{\bar{\phi}^{0}_{v},\phi^{*}_{u_{1}}+\gamma_{v{u_{1}}}\}$ for some $u_{1}\neq v$. If $\bar{\phi}_{v}=\bar{\phi}^{0}_{v}$ then $\bar{\phi}_{v}=\bar{\phi}^{0}_{v}+\gamma_{vv}$ from Lemma 2, and the claim follows. Otherwise, $\bar{\phi}_{v}=\phi^{*}_{u_{1}}+\gamma_{v{u_{1}}}$. Applying this argument repeatedly to components of $\phi^{*}$, it follows that there exist $u_{2},\ldots,u_{k}\in T$, for some $k\leq\left|T-1\right|$, satisfying
$\phi^{*}_{u_{l}}=\phi^{*}_{u_{l+1}}+\gamma_{u_{l}u_{l+1}},$ for $l=1\ldots,k-1,$ and $\phi^{*}_{u_{k}}=\bar{\phi}^{0}_{u_{k}}.$ (7)
Lemma 2 (triangle inequality) implies that $\bar{\phi}_{v}\bar{\phi}^{0}_{u_{k}}+\sum_{l=1}^{k-1}\gamma_{u_{l}u_{l+1}}+\gamma_{vu_{1}}\geq\bar{\phi}^{0}_{u_{k}}+\gamma_{vu_{k}}$. Next, from the feasibility of $\phi^{*}$ and (7), $\bar{\phi}_{v}=\phi^{*}_{v}\leq\phi^{*}_{u_{k}}+\gamma_{vu_{k}}=\bar{\phi}^{0}_{u_{k}}+\gamma_{vu_{k}}$, so it actually holds as an equality. A similar line of reasoning is used to prove the formula for $\underaccent{\bar}{\phi}_{v}$. ∎
The one-dimensional projection established by Proposition 4 is graphically depicted by Figure 2. Note that following Proposition 4, $\mathcal{U}_{\text{SB}}$ can be alternatively written as
$\mathcal{U}_{\text{SB}}=\{\boldsymbol{\phi}\in\mathbb{R}^{|T|}:\underaccent{\bar}{\phi}_{v}\leq\phi_{v}\leq\bar{\phi}_{v},\;\forall v\in T,\;\;\quad|\phi_{v}-\phi_{u}|\leq\gamma_{vu},\;\forall v,u\in T\}.$
Two dimensional examples of $\mathcal{U}_{\text{SB}}$ are graphically illustrated in Figure 3. The two-dimensional case will bear a special importance compactly reformulating the robust problem (2.2). Next, we reformulate the second set of constraints (3d). Again, due to Observation 3, for each $v,u\in T$, the inner maximization is
$\max_{\boldsymbol{\phi}\in\mathcal{U}_{\text{SB}}}\left\{\phi_{v}d_{v}(x)-\mu\phi_{u}d_{u}(x)\right\}=\max_{(\phi_{v},\phi_{u})\in P_{\{v,u\}}(\mathcal{U}_{\text{SB}})}\left\{\phi_{v}d_{v}(x)-\mu\phi_{u}d_{u}(x)\right\}.$ (8)
The next proposition establishes that the projection onto a pair of indices $u,v\in T$, $P_{\{v,u\}}(\mathcal{U}_{\text{SB}})$, can be expressed as a polyhedral set defined by six constraints.
[…]
Figure 3: An illustration of two different two-dimensional projections of $\mathcal{U}_{\text{SB}}$ for voxels $v$ and $u$. The shaded square depicts the $\delta$-bounds on the deviation from the measured $\widehat{\phi}$. The gray $45^{0}$ band depicts the $\gamma_{vu}$ bound on the radiosensitivity difference between the voxels. The green polygons are the projected uncertainty sets. The red points indicate the vertices that potentially maximize (8).
Proposition 5.
[…]
Proof.
Proof. Let $u,v\in T$, and for convenience, define also $\tilde{\mathcal{U}}_{u,v}=\{(\phi_{v},\phi_{u}):-\gamma_{vu}\leq\phi_{v}-\phi_{u}\leq\gamma_{uv}\}$, to show that $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})\subseteq\mathcal{U}_{u,v}$. The contractive property of the projection onto a plane, as a convex set, implies that $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})\subseteq P_{u}(\mathcal{U}_{\text{SB}})\times P_{v}(\mathcal{U}_{\text{SB}})$. Then, by Proposition 4-(i) and the definition of $\mathcal{U}_{\text{SB}}$, it follows that
$P_{\{u,v\}}(\mathcal{U}_{\text{SB}})\subseteq(P_{u}(\mathcal{U}_{\text{SB}})\times P_{v}(\mathcal{U}_{\text{SB}}))\cap\tilde{\mathcal{U}}_{uv}=([\underaccent{\bar}{\phi}_{u},\bar{\phi}_{u}]\times[\underaccent{\bar}{\phi}_{v},\bar{\phi}_{v}])\cap\tilde{\mathcal{U}}_{uv}=\mathcal{U}_{u,v}.$
To show the reverse inclusion, let $(\tilde{\phi}_{u},\tilde{\phi}_{v})\in{\mathcal{U}}_{u,v}$. Consider extending this two dimensional vector to a $\left|T\right|$-dimensional vector in $\mathcal{U}_{\text{SB}}$ in order to prove that $(\tilde{\phi}_{u},\tilde{\phi}_{v})\in P_{\{u,v\}}(\mathcal{U}_{\text{SB}})$. Now, consider some $w\in T\setminus\{u,v\}$ and define $\tilde{\phi}_{w}=\max\{\tilde{\phi}_{v}-\gamma_{vw},\tilde{\phi}_{u}-\gamma_{uw},\underaccent{\bar}{\phi}_{w}\}$. This choice of $\tilde{\phi}_{w}$ satisfies the lower bounds imposed by $\tilde{\phi}_{u},\tilde{\phi}_{v}$ and $\underaccent{\bar}{\phi}_{w}$ in the definition of $\mathcal{U}_{\text{SB}}$. Also, to prove the corresponding upper bounds, first note that
$\tilde{\phi}_{w}\leq\max\{\tilde{\phi}_{v}-\gamma_{vw},\tilde{\phi}_{v}+\gamma_{vu}-\gamma_{uw},\tilde{\phi}_{v}+\gamma_{vw}\}\leq\tilde{\phi}_{v}+\gamma_{vw},$
where the first inequality is due to $(\tilde{\phi}_{u},\tilde{\phi}_{v})\in{\mathcal{U}}_{u,v}$ and $\tilde{\phi}_{v}\geq\underaccent{\bar}{\phi}_{v}\geq\underaccent{\bar}{\phi}_{w}-\gamma_{vw}$ (the latter inequality following from Proposition 4-(ii)) and the second inequality is due to Assumption 1 and Lemma 2 (triangle inequality). Similarly, $\tilde{\phi}_{w}\leq\max\{\tilde{\phi}_{u}+\gamma_{vu}-\gamma_{vw},\tilde{\phi}_{u}-\gamma_{uw},\tilde{\phi}_{u}+\gamma_{uw}\}\leq\tilde{\phi}_{u}+\gamma_{uw}$. Finally, the fact that $(\tilde{\phi}_{u},\tilde{\phi}_{v})\in\mathcal{U}_{u,v}$, together with the definitions of $\underaccent{\bar}{\phi}$ and $\bar{\phi}$, imply that $\tilde{\phi}_{w}=\max\{\tilde{\phi}_{v}-\gamma_{vw},\tilde{\phi}_{u}-\gamma_{uw},\underaccent{\bar}{\phi}_{w}\}\leq\max\{\bar{\phi}_{v}-\gamma_{vw},\bar{\phi}_{u}-\gamma_{uw},\underaccent{\bar}{\phi}_{w}\}\leq\bar{\phi}_{w}$. Next, consider (arbitrary) $w^{\prime}\in T\setminus\{u,v,w\}$ in order to verify that $\tilde{\phi}_{w}$ and similarly defined $\tilde{\phi}_{w^{\prime}}$ together satisfy
$-\gamma_{ww^{\prime}}\leq\tilde{\phi}_{w}-\tilde{\phi}_{w^{\prime}}\leq\gamma_{ww^{\prime}}.$ (9)
Starting with the lower bound,
$\tilde{\phi}_{w}-\tilde{\phi}_{w^{\prime}}$ $=\max\{\tilde{\phi}_{v}-\gamma_{vw},\tilde{\phi}_{u}-\gamma_{uw},\underaccent{\bar}{\phi}_{w}\}-\max\{\tilde{\phi}_{v}-\gamma_{vw^{\prime}},\tilde{\phi}_{u}-\gamma_{uw^{\prime}},\underaccent{\bar}{\phi}_{w^{\prime}}\}$
$\geq\min\{\gamma_{vw^{\prime}}-\gamma_{vw},\gamma_{uw^{\prime}}-\gamma_{uw},\underaccent{\bar}{\phi}_{w}-\underaccent{\bar}{\phi}_{w^{\prime}}\}.$
By definition of $\underaccent{\bar}{\phi}$, $\underaccent{\bar}{\phi}_{w}-\underaccent{\bar}{\phi}_{w^{\prime}}\geq-\gamma_{ww^{\prime}}$. By Lemma 2 (triangle inquality) we also have that $\gamma_{vw^{\prime}}-\gamma_{vw},\gamma_{uw^{\prime}}-\gamma_{uw}\geq-\gamma_{ww^{\prime}}$. Proving the upper bound is analogous (by interchanging the roles of $w$ and $w^{\prime}$ and since $\gamma_{ww^{\prime}}=\gamma_{w^{\prime}w}$). Thus, (9) holds for all $w,w^{\prime}\in T$. Finally, we have established that $\tilde{\phi}\in\mathcal{U}_{\text{SB}}$, and accordingly, it follows that ${\mathcal{U}}_{u,v}\subseteq P_{\{u,v\}}(\mathcal{U}_{\text{SB}})$.∎
Following the results of Propositions  4-5, formulation (2.2) with $\mathcal{U}=\mathcal{U}_{\text{SB}}$ can now be compactly reformulated with a finite number of constraints. In particular, in this reformulation, no variables are added and the number of constraints of this formulation is $2|T|$ times that of the nominal problem (2.1), thereby remaining polynomial in the size of the nominal formulation. Consider the formulation
$\underset{\begin{subarray}{c}{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R}\end{subarray}}{\text{maximum}}\hskip-10.0pt$ $\underaccent{\bar}{d}$ (10b)
subject to $\underaccent{\bar}{\phi}_{v}d_{v}(x)\geq\underaccent{\bar}{d},$ $v\in T$ (10c)
$\bar{\phi}_{v}d_{v}(x)-\mu\max\{\bar{\phi}_{v}-\gamma_{vu},\underaccent{\bar}{\phi}_{u}\}d_{u}(x)\leq 0,\hskip-10.0pt$ $(u,v)\in T^{2}:u\neq v$ (10d)
$\min\{\underaccent{\bar}{\phi}_{u}+\gamma_{vu},\bar{\phi}_{v}\}d_{v}(x)-\mu\underaccent{\bar}{\phi}_{u}d_{u}(x)\leq 0,$ $(u,v)\in T^{2}:u\neq v$ (10e)
$\ d_{v}(x)\leq\bar{d}_{k},$ $k\in[K],v\in H_{k}.$ (10f)
The following theorem establishes the correctness of this formulation, and specifically, demonstrates that its optimal solution set coincides with that of (2.2).
Theorem 6.
[…]
Proof.
Proof. The reformulation of constraints (3c) is a straightforward consequence of (5) together and Proposition 4. Constraints (3d) are reformulated based on (8) to a maximization over $P_{\{v,u\}}(\mathcal{U}_{\text{SB}})$ in place of a maximization over $\mathcal{U}_{\text{SB}}$. Due to Assumption 1, this set is given by the result of Proposition 5. Further, since the objective is linear, and since, for each $u,v\in T$, $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})$ is a bound polyhedral set, then by the fundamental theorem of linear programming, there is an optimal $(\phi_{u},\phi_{v})$ that is an extreme point of the set $P_{\{u,v\}}(\mathcal{U}_{\text{SB}})$. Moreover, since $\phi_{v}$ and $\phi_{u}$ have, respectively, a positive and a negative coefficient in (3d), this implies that the points in $\text{ext}(P_{\{u,v\}}(\mathcal{U}_{\text{SB}}))$ are dominated by the points in
$\mathcal{U}^{*}_{u,v}\equiv\Big{\{}(\max\{\bar{\phi}_{v}-\gamma_{vu},\underaccent{\bar}{\phi}_{u}\},\bar{\phi}_{v}),(\underaccent{\bar}{\phi}_{u},\min\{\underaccent{\bar}{\phi}_{u}+\gamma_{vu},\bar{\phi}_{v}\})\Big{\}}\subset\text{ext}(P_{\{u,v\}}(\mathcal{U}_{\text{SB}}));$
that is, for each pair $v,u\in T$, constraint (3d) is satisfied if and only if its left-hand side is nonpositive for these two points in $\mathcal{U}^{*}_{u,v}$. ∎
In Section 3, we will address the computational aspect of solving (2.2.2), specifically, how to deal with the potentially very large number of homogeneity constraints (10d) and (10e).
3 Solving Large-scale RO
The radiation therapy problems that we consider involve tens of thousands of tumor voxels and millions of healthy tissue voxels. Accordingly, formulation (2.2.2) may have hundreds of millions of constraints. Therefore, it may prove difficult to solve or even store this problem on most computers. In order to address this challenge, we propose an algorithm that dynamically generates constraints of (2.2.2) as needed.
A schematic description is outlined by Algorithm 1.
Input : Initial beamlet configuration $x_{0}$, number of initial constraints for each OAR $n_{0}$, number of homogeneity constraints per iteration $n_{S}$, thresholds $\tau_{\operatorname{Infs},1}>\tau_{\operatorname{Infs},2}$ and $\tau_{z}$
Initialize : $\hat{H}_{1}\subset H_{1},\ldots,\hat{H}_{K}\subset H_{K}$ each with $n_{0}$ constraints violated by $x_{0}$, $S\leftarrow\emptyset$;
1 Phase I: Alternate between iteratively generating at most $n_{H}$ constraints  (10f) to each OAR until a feasibility tolerance $\tau_{\operatorname{Infs},1}$ is achieved and the objective function value does not decrease by more than a factor of $\tau_{z}$, and appending at most $n_{S}$ homogeneity constraints (10d) (10e) to set $S\subseteq T^{2}$. If none of the latter constraints are generated switch to Phase II;
Phase II: The same as Phase I, with feasibility tolerance of $\tau_{\operatorname{Infs},2}<\tau_{\operatorname{Infs},1}$ for constraints  (10f). Exit once no additional constraints can be generated.
Algorithm 1 Constraint generation algorithm sketch
A detailed version of the algorithm in pseudocode is given as Algorithm 3 in Appendix A.1. The algorithm is initialized with beamlet configuration $x_{0}$, typically a part of a previously computed solution or an all-ones vector multiplied by some normalized dose. Based on this initial configuration, each OAR is initialized with the most violated $n_{0}$ constraints. This initialization, and the initial emphasis on generating OAR constraints (10f) in the algorithm, ensures that the optimization problem is bounded from the start or that it becomes bounded after a few iterations in phase I of the algorithm. The algorithm consists of two phases. Each one of them alternates between generation of OAR constraints (10f) and (10d) (10e). The first phase involves an OAR constraint infeasibility tolerance $\tau_{\operatorname{Infs},1}$ that is larger than the tolerance applied in the second phase, $\tau_{\operatorname{Infs},2}$. Intuitively, the larger tolerance in the first phase it intended to ensure both that the subproblems become bounded fairly quickly and to delay generation of OAR constraints that may be inactive at optimality until the homogeneity constraints are nearly satisfied.
Next we consider additional constraints that pertain to healthy organs and we develop an extension of the proposed computational method to address such constraints.
4 Dose-Volume Constraints
In many cases, clinical guidelines stipulate, in addition to or in place of strict bounds on the OAR voxel dose, dose-volume constraints. This refers to the practice of allowing some proportion $\alpha\in(0,1)$, usually less than one half, of the voxels to receive a larger dose than the upper bound $\bar{d}_{K}$ applying to $1-\alpha$ of the voxels. In its most general form, a planning optimization problem could involve several dose-volume constraints for each organ. We develop a method for approximately solving large-scale planning problems in the case of a single dose-volume constraint. The single dose-volume constraint setting is clinically important, as illustrated by the brain-tumor dataset examined in the computational section.
In general, dose-volume constraints can be formulated by introducing additional auxiliary binary variables into formulation (2.2.2), turning it into the mixed-integer program,
$\max_{{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R},y\in\mathbb{R}^{|H_{K}|},w\in\{0,1\}^{\left|H_{K}\right|}}$ $\underline{d}$ (11a)
[…]
$d_{v}(x)\leq\bar{d}_{k}$ $k\in[K-1],v\in H_{k}$ (11c)
$d_{v}(x)-y_{v}\leq\bar{d}_{K}$ (11d)
$y_{v}\leq(\hat{d}_{K}-\bar{d}_{K})w_{v}$ $v\in H_{K}$ (11e)
$\sum_{v\in H_{K}}w_{v}\leq\theta.$ (11f)
Here, without losing generality, it is assumed that the dose-volume constraint applies to the $K$th OAR, auxiliary continuous deviation variables $y_{v}$ are defined for each $v\in H_{K}$, and $\hat{d}_{K}$ is an “absolute” upper bound on the dose of each voxel in $H_{K}$. Setting $\theta=\lfloor\alpha|H_{k}|\rfloor$ in (11f) would ensure compliance with a dose-volume guideline specified as a proportion $\alpha$.
However, solving such a formulation to optimality involves a hard combinatorial problem that is neither theoretically nor practically tractable for the high-dimensional problems considered in the current paper. The standard continuous relaxation of (4) replaces the binary decision variables $w\in\{0,1\}^{\left|H_{K}\right|}$ by continuous ones $w\in[0,1]^{\left|H_{K}\right|}$. Therefore, it is straightforward to see that replacing constraint (11f) with $\sum_{v\in H_{K}}y_{v}\leq\theta\cdot(\hat{d}_{K}-\bar{d}_{K})$, and fixing $w_{v}=1$ in (11e), for all $v\in H_{K}$, maintains the optimality of all solutions that are optimal to the continuous relaxation of (4). Hence, fixing $w_{v}=1$ for all $v\in H_{K}$, for $\Theta=\theta\cdot(\hat{d}_{K}-\bar{d}_{K})$, the formulation
$\max_{{\bf x}\in\mathbb{R}^{n}_{+},\underaccent{\bar}{d}\in\mathbb{R},y\in\mathbb{R}^{|H_{K}|}}$ $\underline{d}$ (12a)
[…]
$\sum_{v\in H_{K}}y_{v}\leq\Theta$ (12c)
$y_{v}\leq\hat{d}_{K}-\bar{d}_{K}$ $v\in H_{K}$ (12d)
is a continuous relaxation of (4). As a continuous relaxation, solutions of this formulation (4) with $\Theta=\theta\cdot(\hat{d}_{K}-\bar{d}_{K})$ may provide bounds but tend to be infeasible for (4). On the other hand, solutions of (4) for sufficiently small values of $\Theta$, for example $\Theta=0$, together with a corresponding indicator variable vector $w$, is feasible for (4). This leads us to consider parametric solutions of (4) and in particular solving an equivalent a penalty-based formulation (similar to a Lagrangian relaxation of (4)) that relaxes constraint (12c), excluding constant terms from the objective. In particular, denote the optimal solution set of this penalty-based problem variant as
[…]
Of course, for every value of $\Theta$ in problem (4), there exists a corresponding value of $\beta$ in problem (13) such that the optimal values of $\underaccent{\bar}{d}$ for the two problems coincide. For every $\beta\geq 0$, the notations $s(\beta)\equiv(x(\beta),\underaccent{\bar}{d}(\beta),y(\beta))$ and $z(\beta)$ are used, respectively, to denote an optimal solution (an arbitrary element of the optimal solution set $\mathcal{S}(\beta$)) and the optimal objective value of (13) with penalty parameter $\beta$. The next lemma establishes that $\underaccent{\bar}{d}(\beta)$ and $\sum_{v\in H_{K}}y(\beta)_{v}$ are monotonically non-increasing in $\beta$ (see proof in Appendix B.1).
Lemma 7.
For every $0\leq\beta_{1}<\beta_{2}<\beta_{u}$, every pair of optimal solutions $(x,\underline{d},y)\in\mathcal{S}(\beta_{1})$ and $(\tilde{x},\tilde{\underline{d}},\tilde{y})\in\mathcal{S}(\beta_{2})$ satisfies
$\sum_{v\in H_{K}}\tilde{y}_{v}\leq\sum_{v\in H_{K}}y_{v},$ $\tilde{\underaccent{\bar}{d}}\leq\underaccent{\bar}{d},$ and $\tilde{\underaccent{\bar}{d}}-\beta_{2}\sum_{v\in H_{K}}\tilde{y}_{v}\leq\underaccent{\bar}{d}-\beta_{1}\sum_{v\in H_{K}}y_{v}.$
Lemma 7, together with the fact that there exists a value of $\beta>0$ for which $\sum_{v\in H_{K}}y_{v}(\beta)=0$, implies that for every value of $\Theta\geq 0$, it is possible to find a minimal value of $\beta$ such that (12c) is satisfied. Motivating our method is both monotonicity of $\left|\left|y(\beta)\right|\right|_{1}=\sum_{v\in H_{K}}y(\beta)_{v}$, following Lemma 7, and that the one-norm is a convex approximation of the zero-“norm” ($\left|\left|y(\beta)\right|\right|_{0}$) , which is the number of nonzero components of the deviation variable vector $y$. The one-norm is effectively used to promote sparsity in large-scale machine learning and compressed sensing applications (see, for example, Donoho and Elad (2003)). We expect that solving (13) with appropriate $\beta$ will approximate the problem with dose-volume constraints (11f) reasonably well. Our goal is to find the minimal value of $\beta$ for which the optimal solution of (13) satisfies (11f). To achieve this , LP (13) is solved parametrically, starting from a lower bound on $\beta$ and traversing optimal bases for the given subsets of constraints defining the reduced problem. This algorithm is presented as Algorithm 2.
An additional difficulty in implementing our algorithm lies in the fact that, in practice, (13) is often too large to solve with all constraints present, and the constraints are generated on the fly as in Algorithm 1. Consequently, some of the constraints and variables may be absent from the formulation, and accordingly the reduced cost of these variables, as well as the full basisare not readily available. However, both the reduced cost and full basis are needed for determining the minimal increase in $\beta$ for which the current optimal basis will no longer be optimal, and consequently $\left|\left|y\right|\right|_{0}$ may change. Our parametric algorithm takes this into account by possibly retracting and allowing the value of the parameter $\beta$ to decrease if the computed solution is not adjacent to the preceding basic feasible solution in the polyhedron that incorporates the newly generated constraints.
Input : Initial bounds $0\leq\beta_{l}<\beta_{u}$, BFS optimal for lower bound $\beta=\beta_{l}$, initial constraint subsets $(S,\hat{H}_{1},\ldots,\hat{H}_{K})$
1 while $\left|\left|y_{\hat{H}_{K}}\right|\right|_{0}>\theta$ or current BFS is not adjacent to the previous one in the full polyhedron with constraints $(T,H_{1},\ldots,H_{K})$ do
2       If the current BFS satisfies (or violates) $\left|\left|y_{\hat{H}_{K}}\right|\right|_{0}>\theta$ then minimally increase (respectively decrease) $\beta$ so that current BFS is no longer optimal;
3       Solve restricted LP $(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta)$ to obtain optimal BFS with updated constraint sets $(S,\hat{H}_{1},\ldots,\hat{H}_{K})$
4 end while
Output : $\beta$ and corresponding BFS.
Algorithm 2 Parametric penalty algorithm sketch
A detailed version of this algorithm is given in Appendix A.3. The algorithm applies row and column generation in Step 2 in order to solve $\Phi(S,\hat{H},\ldots,\hat{H}_{K},\beta)$; the details of this procedure are presented as Algorithm 4 in Appendix A.2. This algorithm is a subroutine for solving (13) that is similar to Algorithm 1 in that it solves the problem for a subset of the constraints (rows), but it also generates columns corresponding to $y$ variables as needed (only those components of $y$ corresponding to constraints (11d) and (11e) that have already been generated).
Let $\beta^{*}=\min\{\beta:\exists(x,\underaccent{\bar}{d},y)\in\mathcal{S}(\beta),\|y\|_{0}\leq\lfloor\alpha|H_{K}|\rfloor\}$; that is, $\beta^{*}$ is the minimal penalty parameter for which the optimal solution to (13) satisfies  (11f). Bounds on $\beta^{*}$ can be established using Lagrange multipliers of constraint (12c) for solutions that are optimal to (4) for different values of $\Theta$, specifically, $\beta_{l}$ and $\beta_{u}$ for $\Theta=\lfloor\alpha|H_{K}|\rfloor(\hat{d}_{K}-\bar{d}_{K})+\epsilon$ (for some $\epsilon>0$) and $\Theta=0$, respectively. These bounds are proven in Appendix B.2. The following proposition establishes the correctness of Algorithm 2 given such bounds.
Proposition 8.
Suppose that LP (13) has a unique optimal basis for all $\beta\in[\beta_{l},\beta_{u}]$. Then, Algorithm 2 determines $\beta^{*}$ in a finite number of steps.
The proof of the proposition is deferred to Appendix B.3.
5 Experiments
In our experiments, we focus on a particular case study based on real brain data of Patient 4 from the TCIA archive (Clark et al. 2013), referred to as the Brain dataset in Table 1. Due to the very large number of constraints in (2.1), and especially in (2.2.2), for this dataset, we also use a smaller liver imaging dataset (Craft et al. 2014) to evaluate the computational performance of our dynamic generation of OAR and homogeneity constraints in Algorithm 3. FMISO-PET images are not available for this smaller dataset, so the radiosensitivity data are created synthetically using the following formula for $v\in T$: $\phi_{v}=0.85+0.15\times\frac{\max_{u\in T}\{f_{u}\}-f_{v}}{\max_{u\in T}\{f_{u}\}-\min_{u\in T}\{f_{u}\}}$, where vector $f$ follows a multivariate normal density over the PTV voxels, with the mean set as the average voxel x-y-z coordinates, and the covariance proportional to the covariance of the x-y-z coordinates (scaled by 0.2). Details of both datasets are given in Table 1.
Table 1: Imaging data summary
Dataset PTV OAR
$\left|T\right|$ $k$ Name $\left|H_{k}\right|$ $\bar{d}_{k}$ bounds [^1] Based on (ICR 2010)
Brain (Patient 4) 32,367 1 Exterior 1,870,179 100
2 Brain 1,360,182 62
3 Chiasm 2,003 54
Liver 6,954 1 Liver 77,657 60
2 Heart 32,983 50
3 Entrance 146,462 60
5.1 Computational Performance
Our implementation uses the Julia programming language, version 1.6.2, the JuMP package (Lubin and Dunning 2015), and the Gurobi solver, version 9.1.2. The experiments are run on 40 core Linux servers each equipped with Intel Xeon E5-2650, 25 MB cache, 2.3 GHz base frequency CPUs, and 128 GB of RAM. Algorithm 1 (appearing in greater detail as Algorithm 3 in Appendix A.1) is invoked for solving the robust model (2.2.2) and the nominal model (2.1), with parameter values $\tau_{\operatorname{Infs},1}=10$, $\tau_{\operatorname{Infs},2}=0$, $\tau_{z}=10^{-2}$, and $n_{H}=400$. The homogeneity parameter is set to $\mu=1.15$ for Liver instances and $\mu=1.1250$ for Brain instances for the nominal experiments, and $\mu=1.1875$, which obtains a similar homogeneity and biologically-adjusted dose in the nominal scenario. Table 2 shows the running times achieved when solving the nominal model (2.1), which involves only $\left|T\right|$ homogeneity constraints. The results show that dynamic constraint generation is not needed to solve the nominal model but with careful tuning of the constraint generation algorithm, it can improve on the running time required to solve the full model (also yielding significant savings in terms of memory consumption), especially for larger data instances such as Brain. Table 3 shows the running times when applying the robust model to both datasets. The parameters of the uncertainty set are set to $\delta=0.04$ and $\gamma=0.04$ for Liver, and $\delta=0.08$ and $\gamma=0.04$ for Brain. Here, it is evident that dynamic constraint generation is necessary in order to be able to solve both models when the number of homogeneity constraints is $O(\left|T\right|^{2})$. However, dynamically generating all violated constraints may result in excessive running times. In the example shown in Table 3, dynamically generating all violated constraints in each iteration results in a running time of nearly four hours for the Brain instance. In preliminary experiments, with a different choice of parameters than experimented with here, when generating all violated constraints, we have observed cases where the running times took more than 48 hours to solve, and cases memory consumption exceeded its limits.
Table 2: Running time in seconds and number of constraints generated in solving the nominal model (2.1). Here, constraint generation, if indicated, implies that all violated constraints are generated in each iteration.
Hom cons OAR cons Liver Brain
generation initial - $n_{0}$ Time # generated Time # generated
No 2,000 126.3 19,908 586.7 70,737
No 8,000 82.3 37,908 731.7 82,737
No all 453.4 271,010 5207.2 3,297,098
Yes 2,000 189.5 22,858 3,235.1 72,160
Yes 8,000 175.9 37,531 3,458.7 77,730
Yes all 563.7 265,977 5,117.7 3,264,928
Table 3: Running time in seconds and number of constraints generated in solving (2.2.2) using Algorithm 3.
Hom cons OAR cons Liver Brain
per iter - $n_{s}$ initial - $n_{0}$ Time # generated Time # generated
2,000 2,000 3,718.6 92,890 4,940.4 82,160
2,000 all 10,215.3 335,528 12,861.4 3,271,931
8,000 2,000 3,857.8 130,343 5,255.7 106,261
8,000 all 10,033.0 385,506 12,029.5 3,286,449
all all 8,133.5 408,747 14,356.1 3,424,511
5.2 Robust Model Evaluation
Next, the spatially bound uncertainty set $(\mathcal{U}_{\text{SB}})$, introduced in Section 2.2, will be constructed and customized for our patient data. Subsequently, the benefits of both the robust model (2.2.2) based on this uncertainty set, and the dose-volume methodology of Section 4, will be examined and discussed.
5.2.1 Uncertainty Set Construction
The method of constructing our uncertainty set is justified in Section 2.2 using the data of brain cancer Patient 1 from the TCIA archive in Clark et al. (2013), illustrated in Figure 1. The graphs in Figure 4, which are based on the Brain data of Table 1 (Patient 4 from the archive), demonstrate a similar pattern to the one that emerges from Figure 1. Both figures clearly illustrate that the radiosensitivity distance increases in the geometric voxel distance, up to a geometric distance of approximately ten. To estimate the difference in radiosensitivity between voxel pairs as a function of the voxel distance $\Delta$, we adopt the following model with parameters $\alpha_{0}$, $\alpha_{1}$, $\alpha_{2}$, and $\gamma$ (initially 0),
$\Gamma(\Delta)=\begin{cases}\gamma+\alpha_{0}+\alpha_{1}\Delta+\alpha_{2}\log(\Delta)&\Delta\leq 10\\ \Gamma(10)&\text{otherwise.}\end{cases}$
The estimated $\Gamma$ is personalized to Patient 4’s PET data. It is estimated through a constrained least squares fit of the radiosensitivity difference $p$-th percentile curve (see Figure 4), for voxel distances up to $\Delta=10$. The least squares problem solved, is constrained in our case, so that the fitted curve upper bounds the 98th percentile graph. Having obtained the best fit curve, the constant $\alpha_{0}$ is then slightly increased so that the curve upper bounds the same $p$-th percentile for all $\Delta\geq 1$. The constant $\gamma\geq 0$ is used to adjust the model to be even more conservative. The final estimated parameter values are $\alpha_{0}=0.0292761$, $\alpha_{1}=-0.0013514$, and $\alpha_{2}=0.0128265$.
Figure 4: Radiosensitivity difference vs. voxel distance for PTV voxel pairs of Patient 4 in the TCIA archive Brain dataset.
Evidently, $\Gamma$ satisfies Assumption 1 for all $\gamma\geq 0$. Moreover, for $\gamma=0$, by construction, $\Gamma$ upper bounds the $98$th percentile graph, and for $\gamma=0.04$, $\Gamma$ upper bounds the graph of the maximum, as seen in Figure 4(A).
The magnitude of the uncertainty set parameter $\delta$, which bounds the deviation of a voxel’s radiosensitivity from its calculated value can be estimated by a sensitivity analysis of underlying theoretical equations in the uncertain constants; see for example the discussion in Appendix C.
5.2.2 Robust Formulation Experiments
We now evaluate optimal solutions of spatially robust model (2.2.2) and compare them with optimal solutions of the nominal formulation (2.1) for the Brain dataset. We also compare with the robust box model that is a robust model (2.2) under a box uncertainty set ($\mathcal{U}_{B}$). The solutions to the robust box model can also be obtained by solving (2.2.2) with $\underaccent{\bar}{\phi}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\underaccent{\bar}{\phi}^{0}}$ and $\bar{\phi}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\bar{\phi}^{0}}$ by setting $\gamma_{uv}$ in the $\mathcal{U}_{\text{SB}}$ uncertainty set to a large value (it suffices to set $\gamma_{uv}=1$) for all $u,v\in T$. Thus, the robust box case can be viewed as a special (simple) case of our suggested approach.
Optimal solutions of (2.2.2) (with $\mathcal{U}_{\text{SB}}$) are computed for a range of uncertainty set parameter values, $\delta=0.02,0.04,\ldots,0.14$ and $\gamma=0.01,\ldots,0.05$, and a range of homogeneity parameter values, $\mu=1.0625,1.0750,\ldots,1.4000$. Note that for some combinations of uncertainty set and homogeneity parameters, formulation (2.2.2) has $x=0$ as the only feasible solution; such solutions are omitted from the graphs. For the box uncertainty set ($\mathcal{U}_{B}$), the solution were computed with the same values for $\delta$, however, a larger range of homogeneity parameters, $\mu=1.0625,1.0750,\ldots,1.4750$, is explored to be able to obtain nonzero solutions for all $\delta$ values. Moreover, in order to obtain a more accurate value of $\mu$ below which there is no nonzero dose solution, for each $(\gamma,\delta)$ pair, the $\mu$ values between the first nonzero dose solution and the last zero dose solution were refined, first in increments of $0.0025$, then in incremenets of $0.0005$, and finaly in increments of $0.0001$. To evaluate a plan involving beamlet decision variable vector $x^{*}$ under uncertainty set $\mathcal{U}$, which could be a singleton when considering the nominal case, empirical measures of worst-case dose and homogeneity are defined as follows
$\hat{\underaccent{\bar}{d}}=\hat{\underaccent{\bar}{d}}(x^{*},\mathcal{U})=\min_{\hat{\phi}\in\mathcal{U}}\min_{v\in T}\hat{\phi}_{v}d_{v}(x^{*})$ and $\hat{\mu}=\hat{\mu}(x^{*},\mathcal{U})=\max_{\hat{\phi}\in\mathcal{U}}\max_{u,v\in T}\frac{\hat{\phi}_{u}d_{u}(x^{*})}{\hat{\phi}_{v}d_{v}(x^{*})};$
$\hat{\underaccent{\bar}{d}}$ and $\hat{\mu}$ are used as abbreviations when $x^{*}$ and $\mathcal{U}$ are unambiguous in the given context.
Before comparing the performance of solutions that are optimal to (2.2.2) under different scenarios, it should be noted that the optimal solution value of (2.2.2) decreases in the uncertainty set parameters $\delta$ and $\gamma$ (see Figure 13 in Appendix D). Note that combinations of $\mu$, $\delta$, and $\gamma$ resulting in a zero beamlet intensity configuration ($x={\mathbf{0}}$) are omitted from all figures and are interpreted as “infeasible” plans.
Figure 5: Nominal biologically-adjusted performance of solutions optimal to model (2.2.2) for $\delta=0.14$ and different values of $\gamma$, compared to the nominal solution.
A “cost-of-robustness” can be measured by evaluating a robust solution in the nominal scenario where there is no uncertainty. Figures 5 and 6 illustrate this phenomenon. The nominal formulation (2.1) corresponds to setting $\delta=0$ and $\gamma$ sufficiently large in $\mathcal{U}_{\text{SB}}$; for the particular dataset being considered (Patient 4), the nominal formulation is equivalent to a setting of $\delta=0$ and $\gamma\geq 0.04$ in formulation (2.2.2)). On the other hand, uncertainty sets with $\gamma\leq 0.03$ exclude the nominal scenario. Consequently, models with smaller $\gamma$ values tend to perform poorly, at least in terms of the minimum biological dose, in the nominal scenario. Figure 5 shows the performance in the nominal scenario, for $\delta=0.14$, with different values of $\gamma$ and $\mu$ in (2.2.2) (a curve is presented for each value of $\gamma$, obtained by varying the values of $\mu$). In general, the “cost of robustness” appears more evident as $\gamma$ decreases: the minimal $\hat{\mu}$ that is attained with a nonzero dose solution increases, and $\hat{\underaccent{\bar}{d}}$ somewhat decreases. This can be explained by the fact that the nominal scenario is excluded from the uncertainty set when $\gamma\leq 0.03$, and the distance between this scenario and the uncertainty set increases as $\gamma$ decreases. The figure also shows that the robust box solution and the robust spatially bound solution with $\gamma=0.04$ exhibit fairly similar performance in the nominal scenario but the robust box appears less stable and demonstrates a notable drop in dose for biological homogeneity values slightly larger than 1.1. Note that a lack of stability stems from the fact that the measured $\hat{\mu}$ and $\hat{\underaccent{\bar}{d}}$ values typically differ from the values of $\mu$ and $\hat{\underaccent{\bar}{d}}$, respectively, that are used in optimizing (2.2.2). Consequently, even slight changes in the parameter $\mu$ may in some cases induce a different optimal solution that is associated with significantly different values of the measured $\hat{\mu}$ and $\hat{\underaccent{\bar}{d}}$. In this case, when increasing $\mu$ from $1.3603$ to $1.3607$, the nominal performance of the robust box model solution improves with respect to both homogeneity and dose, from $\hat{\mu}=1.105$ and $\hat{\underaccent{\bar}{d}}=37.4$ to $\hat{\underaccent{\bar}{\mu}}=1.1$ and $\hat{\underaccent{\bar}{d}}=55.1$.
Figure 6 shows nominal performance curves for $\gamma=0.02,0.04,1$, $\delta=0.02,0.04,\ldots,0.1{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}$, and a range of $\mu$ values (each curve is obtained by varying the value of $\mu$). For each of the $\gamma$ values, the “cost of robustness” becomes more evident as $\delta$ increases: the curves for larger $\delta$ values appear to be nearly dominated in the $\hat{\mu}-\hat{\underaccent{\bar}{d}}$ space by the curves obtained for smaller $\delta$ values. The robust box results appear much less stable than the corresponding SR model results for $\gamma\in\{0.02,0.04\}$, especially for values of $\hat{\mu}\leq 1.11$, with doses 25Gy lower than the optimal nominal solution compared to at most 5Gy lower for $\gamma=0.02$ and $0.04$. Recall that solutions of (2.2.2) with $\gamma=0.02$ are not designed to protect against the nominal realization (which is excluded from the uncertainty sets with $\gamma<0.04$). Hence, the solutions obtained in the case where $\gamma=0.02$ have a biologically-adjusted dose of approximately $0.81$ Gy less than the corresponding optimal nominal dose for all $\hat{\mu}>1.11$. When increasing $\gamma$ to $0.04$, the uncertainty set captures the nominal solution and so the dose obtained for $\hat{\mu}\geq 1.11$ become equal to those of the nominal model, albeit the minimal $\hat{\mu}$ for which there was a nonzero dose is higher than the nominal model, even for small values of $\delta$. This trend persists when moving to the robust box model, which results in almost optimal nominal performance for $\delta\leq 0.04$.
A $\hat{\underaccent{\bar}{d}}$ vs. $\hat{\mu}$, $\gamma=0.02$
B $\hat{\underaccent{\bar}{d}}$ vs. $\hat{\mu}$, $\gamma=0.04$
C $\hat{\underaccent{\bar}{d}}$ vs. $\hat{\mu}$, robust box ($\gamma=1$)
Figure 6: Nominal biologically-adjusted performance of solutions optimal to models (2.1) and (2.2.2). (for $\gamma=0.02$, $\delta=0$ is infeasible.)
Next, we consider the physical dose performance of the obtained solutions. Although physical performance measures do not account for the variable radiosensitivity, and planning based on such measures may not utilize the information that has been recently made available by advanced biomarkers such as FMISO, they are still important as the ongoing and current basis for treatment guidelines and clinical practice. Keeping in mind that physical performance is not explicitly maximized in the optimization formulations (2.1) or (2.2.2), it may be interesting to examine how the resulting planning solutions fare with respect to the physical performance measures. Figure 7 compares solutions of (2.2.2) with $\delta=0.14$ and different values of $\gamma$, the robust box solution (also with $\delta=0.14$), and the nominal solution. Figure 6A depicts the physical homogeneity vs. minimal physical dose of the different models, and Figure 6B shows the EUD (equivalent uniform dose). The EUD measure was introduced as a single number representing the dose throughout the PTV (Niemierko 1997, Choi and Deasy 2002), and is given by the formula $EUD=\left(\frac{1}{|T|}\sum_{v\in T}d_{v}^{\alpha}\right)^{\frac{1}{\alpha}},$ where, as in Nohadani and Roy (2017) we use $\alpha=-10$. Surprisingly, both the optimal nominal solution and the robust box solution exhibit poor performance compared with the solutions of SR model (2.2.2) with respect to both physical homogeneity and physical minimal PTV dose, for physical $\mu$ values in the interval $(1.07,1.10)$. Specifically, the solutions of the nominal and robust box models are less stable and, with both physical dose and EUD up to 20Gy lower than for $gamma\leq 0.05$, and a minimal physical homogeneity value of around $1.07$ compares to $1.02$ for $\gamma\leq 0.05$. Similar behavior can be observed for other values of $\delta$ tested; see Figure 15 in Appendix D. A possible explanation for this behavior is that under both the nominal and the box uncertainty plans, the dose is allocated in a way that is focused mainly on the PTV voxels with the lowest radiosensitivities (following the max-min formulation). The spatially bound uncertainty may smooth out the PTV radiosensitivity uncertainty, resulting in a worst case that considers a greater number of voxels, including voxels for which the radiosensitivity is nominally higher, thus increasing the minimal physical dose.
A Minimum physical dose
B EUD
Figure 7: Physical performance of solutions optimal to (2.2.2) with $\delta=0.14$ and a range of $\gamma$ values, compared to the nominal solution.
A FMISO PET
B Nominal
C Spatial
D Box
Figure 8: (7A) – PET, pink and red levels indicating low oxygen. (7B)-(7D) – physical dose.
Figure 9: Comulative dose-volume histograms.
Figures 8-9 display the physical dose obtained by the nominal solution, the robust box solution with $\delta=0.14$, and the robust spatial solution of (2.2.2) with $(\delta,\gamma)=(0.14,0.04)$. For each model, the value of the homogeneity parameter $\mu$ is chosen so that it yields the highest minimal physical dose while maintaining physical homogeneity less than 1.1 (which is a standard clinical requirement). All three dose images in Figure 7B-7D are aligned with the same PET/CT FMISO scan slice 7A. The robust box and nominal solutions apply wider physical dose ranges to the PTV, 61.5-67.5 Gy for the robust box and 62-68 Gy for the nominal, compared to the robust spatial solution, with a range of 63-65 Gy. Furthermore, the nominal and robust box solutions create dose “hot spots” of approximately 100 Gy in the perimeter of the exterior OAR, while the robust spatial model solution applies a maximal physical dose of less than 71 Gy to this OAR. This may be due to the fact that the nominal and robust box solutions use fewer beams, with some of the beams set to a higher intensity, than the robust spatial solution. Thus, while the robust spatial model retains approximately the same adjusted dose as the nominal solution, it accounts for the variability and uncertainty in the radiosensitivities. It also appears to avoid excessively high OAR doses.
A “True model” with $\gamma=0.02$
B “True model” with $\gamma=0.04$
Figure 10: Worst-case biologically-adjusted performance ($\hat{\mu}$, $\hat{\underaccent{\bar}{d}}$) of solutions optimal to model (2.2.2) with misspecified $\gamma$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\delta=0.1$ and $\gamma$ as indicated. The performance of the nominal solution is also shown for comparison.
Finally, to evaluate the cost of misspecifying the uncertainty set parameters, Figures 10 and 11 depict the worst-case performance of the various solutions under the assumption that the correct uncertainty models differ with respect to the parameter values assumed in formulation (2.2.2). In Figure 10 the robust solutions use the “true value” of $\delta=0.1$, so the misspecification is only in terms of $\gamma$. Specifically, $\gamma=0.02$ and $\gamma=0.04$ in Figure 9A and Figure 9B, respectively. Note that while the worst-case performance deteriorates as the value of $\gamma$ used to derive the solution deviates from the true value, the dose declines by at most (and is generally much less than) $1.1$ Gy. Importantly, the spatially robust models may achieve homogeneity values that are close to the homogeneity of the correct model, improving the homogeneity by more than $0.05$ relative to the nominal and robust box solutions. This is evident for most values of $\delta$ (see Figures 16 and 17 in Appendix D). Moreover, although the robust box and nominal model appear to attain a similar biologically-adjusted dose as the true model for high values of $\hat{\mu}$, the performance of these models appears much less stable for desirable (lower) values of $\hat{\mu}$. Notably, the robust box obtains doses significantly lower than all models considered including the nominal one for some values of $\hat{\mu}$. Hence, even when $\gamma$ is misspecified in (2.2.2), it may still yield nearly the same homogeneity, $\hat{\mu}$, perhaps at a slight cost to $\hat{\underaccent{\bar}{d}}$.
A “True model” with $\delta=0.02$
B “True model” with $\delta=0.14$
Figure 11: Worst-case biologically-adjusted performance ($\hat{\mu}$, $\hat{\underaccent{\bar}{d}}$) of solutions optimal to (2.2.2) with misspecified $\delta$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\gamma=0.04$ and $\delta$ as indicated
Figure 11 depicts experimental results for evaluating the effect of misspecifying $\delta$ in the uncertainty model. In this set of experiments, the true uncertainty set, as well as the uncertainty sets assumed in solving (2.2.2), all share the same value of $\gamma=0.04$, but differ in the values of $\delta$. While the misspecification of $\delta$ seems to cause only a modest decrease in the value of $\hat{\underaccent{\bar}{d}}$ (of up to $0.5$ Gy) compared with the value obtained under the true $\delta$, while the homogeneity, $\hat{\mu}$, appears to be more significantly affected. Specifically, underestimating $\delta$ increases $\hat{\mu}$ of up to $0.03$ (for a similar dose), while overestimating $\delta$ may lead to little or no improvement in the homogeneity (a small decrease in $\hat{\mu}$). Further, the $\hat{\mu}$ values attained for the nominal model are significantly worse (larger than the spatially robust models by up to 0.06) for the higher true value of $\delta=0.14$, compared to being approximately 0.03 larger than all spatially robust models when the true $\delta=0.02$.
5.3 Dose-Volume Experiments
In this section, we conduct experiments using Algorithm 2, introduced in Section 4 for handling dose-volume constraints. The experiments in this segments were run using Julia 1.3.0, and Gurobi solver 9.0.0, on an Intel Xeon Gold 6254 CPU with a 3.10GHz base frequency, 24.75 MB cache, and 376 RAM.
In order to speed up the computation, we may first narrow down the search interval further by performing a preliminary bisection search, and then proceed by running Algorithm 5 given this reduced interval. Note that invoking the bisection search may sacrifice some of the theoretical properties guaranteed by the initial lower bound, thus serving only as an effective heuristic. Further, in order to deal with possible degeneracy in Algorithm 2 we determine adjacency of basic feasible solutions only based on the $x$ and $y$ variables. In these experiments, we focus on the dose-volume constraint for the brain (OAR 2 of Patient 4). Specifically, following the ICRU guidelines (ICR 2010), $99\%$ of the brain should receive less than $\bar{d}_{2}=60$ Gy while $\hat{d}_{2}=62$ Gy; that is, $\alpha=0.01$ in constraint (11f). For these experiments we consider both the nominal and spatially robust settings with the same parameters as in Section 5.1.Recall that Algorithm 5 determines the minimal $\beta$ value for which the solution satisfies this constraint. Intuitively, a small $\beta$ and corresponding penalty term in the objective should correspond to a small change in the objective compared with the non-penalized problem (where $\beta=0$). Thus, the constraint is expected to be satisfied with a slight decline in the optimal objective value relative to that achieved when $\beta=0$.
Figure 12 depicts the value of the adjusted dose $\underaccent{\bar}{d}$, and the $\ell_{1}$ and $\ell_{0}$ norms of the deviation vector $y$, as a function of $\beta$, for the different values of $\beta$ obtained by the algorithm for both cases. In order to run the algorithm, we first computed $\bar{\beta}\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.22}\times 10^{-2}$ for the nominal case, and $\bar{\beta}\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.58}\times 10^{-2}$ for the specially robust case. We used $\underaccent{\bar}{\beta}=0$ in both cases, while the true lower bounds are $\underaccent{\bar}{\beta}\approx 1.73\times 10^{-6}$ for the nominal setting and $\underaccent{\bar}{\beta}\approx 1.36\times 10^{-6}$ for the robust setting. The algorithm determined that $\beta^{*}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.72}\times 10^{-5}$ and $\beta^{*}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.4}\times 10^{-5}$ are the smallest $\beta$ values that satisfy $\|y\|_{0}\leq\lfloor\alpha|H_{2}|\rfloor=13,601$ for the nominal and the spatially robust case, respectively (thus satisfying the dose-volume constraint).
A $\underaccent{\bar}{d}$ vs. $\beta$
B $\left|\left|y\right|\right|_{1}$ and $\left|\left|y\right|\right|_{0}/\left|H_{2}\right|$ vs. $\beta$
Figure 12: Minimum biologically-adjusted dose $\underline{d}$ and deviation vector $y$ norms vs. $\beta$ values enumerated by Algorithm 2 for the nominal and spatially robust settings.
We compared the solutions obtain by Algorithm 5 in both of the spatially robust and nominal settings, with a solution obtained from problem (2.2) with an additional CVAR constraint, as suggested by Romeijn et al. (2003). Table 4 shows these results, comparing also with the unresricted case, where the dose-volume constraints are not imposed (having $\beta=0$ in (13)).
Table 4: Comparison of Algorithm 5 with adding a CVaR constraint
Setting Method $\underaccent{\bar}{d}$ EUD % Running
Violation Time (sec)
Nominal Unrestricted 57.26 64.85 3.91% 2904
Algorithm 5 56.92 64.50 1.00% 55784
CVaR 56.46 64.03 0.41% 87192
Robust Unrestricted 52.08 64.73 3.89% 4327
Algorithm 5 51.81 64.33 1.00% 68980
CVaR 51.38 63.80 0.41% 105366
Algorithm 5 only reduces the dose by $0.59\%$ and $0.52\%$ compared to the unrestricted case, in the nominal and robust settings, respectively. Also, note that CVaR results in lower values of both minimal biologically-adjusted dose and EUD, compared with Algorithm 5. This is especially evident in the nominal setting. This is due to the CVaR constraint being a safe approximation of the dose-volume constraint, resulting in more conservative solutions to the planning problem. Indeed, the percent of brain voxels exceeding the $\bar{d}_{2}=60$Gy bound is approximately $0.4\%$ using the CVaR approach, which is lower than the required $1\%$. Finally, the CVaR constraint requires a variable for each of the brain voxels, which may be computationally intensive; without initializing to the unrestricted solution, we found that CVaR would not solve even after 72 hours. With this particular initialization, we are able to solve for CVaR but it takes about 1.5 times as long as Algorithm 5, as displayed in Table 4. Thus, when Algorithm 5 can be applied, it provides an alternative to the CVaR approach that is both superior in terms of performance and in terms of the running time.
6 Conclusions and Future Work
We proposed a novel spatially based uncertainty set to model uncertainty in tumor radiosensitivity of radiation treatment based on measured biomarkers. Naively incorporating this uncertainty set in robust voxel-based FMO models would lead to an unmanageable increase in the already large problem size. Alternatively, we compactly reformulate the robust problem by proving that at most a few uncertainty set extreme points need to be considered for each constraint. While theoretically tractable, the new compact formulation may still contain over a billion constraints. We first show that although the compact formulation cannot be solved directly it can be solved within a reasonable running time by carefully generating constraints as needed. Extensive experiments demonstrate advantages of the proposed uncertainty set in terms of the adjusted homogeneity-dose tradeoff over the nominal and robust box-uncertain models.
This work has focused on spatial aspects of biomarker related uncertainty in a voxel-level planning model (also known as “dose painting”). In future work, we would like to extend the proposed model to include temporal aspects, taking into account fractionation of treatments as well as a possible additional image scan midtreatment. This leads to a challenging two-stage adaptive robust problem, where the spatial uncertainty changes over time. Finally, it is worth noting that the study of robust and adaptive planning under the condition that neighboring elements uncertainty is connected has application beyond radiation therapy. Communicable disease, traffic congestion, and even agriculture are all examples of situations where uncertainty among neighbors in space and time is connected.
Data Availability Statement
All data used in this paper is based on data available from public repositories whose references are indicated and with processing as described in Section 5. This data in its final processed form will be made available by the authors upon reasonable request.
Appendix A Detailed Algorithm Descriptions
A.1 Row Generation Algorithm
[…]
Input : Initial beamlet configuration $x_{0}$, number of initial constraints for each OAR $n_{0}$, number of homogeneity constraints per iteration $n_{S}$, number of OAR constraints per iteration $n_{H}$, thresholds $\tau_{\operatorname{Infs},1}>\tau_{\operatorname{Infs},2}\geq 0,\tau_{\operatorname{Infs},3}\geq 0,\tau_{z}>0$
Initialization : $S\leftarrow\emptyset$, $\hat{H}_{1},\ldots,\hat{H}_{K}$ based on $x_{0}$ with $n_{0}$ constraints in each OAR, $z_{\text{prev}}\leftarrow\infty$, $m_{S}\leftarrow 1$, $m_{H_{k}}\leftarrow 0,\;\forall k$;
1 while $\sum_{k=1}^{K}m_{H_{k}}+m_{S}>0$ do
2       if $m_{S}>0$ then
3            Solve $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K})$ and obtain its optimal solution $s$ and optimal value $z$;
4       end if
[…]
6             $z_{\text{prev}}\leftarrow z$;
7             for $k=1,\ldots,K$ do
[…]
9             end for
10            if $\sum_{k=1}^{K}m_{H_{k}}>0$ then
11                   Solve $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K})$ and obtain its optimal solution $s$ and optimal value $z$;
12             end if
13
14       end while
[…]
16
17 end while
Output : $s$, $z$
Algorithm 3 Row generation algorithm
A.2 Row and Column Generation Algorithm
Here, $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta)$ denotes an instance of problem (13) with constraints (10d) and (10e) defined only for a subset $S\subseteq T^{2}$, constraints (11c) defined only for subsets $\hat{H}_{1}\subseteq H_{1},\ldots,\hat{H}_{K-1}\subseteq H_{K-1}$, and constraints (11d),(11e) defined for subset $\hat{H}_{K}\subseteq H_{K}$.
Accordingly, solutions of this problem are of the form $s=(x,\underaccent{\bar}{d},y_{\hat{H}_{K}})$. For solutions that are optimal to this problem, note that, for all $v\in\hat{H}_{K}$, $y_{v}=\max\{d_{v}(x)-\bar{d}_{K},0\}$. For all $v\in H_{K}\setminus\hat{H}_{K}$, we similarly define, for the given solution $s$, $y_{v}=\max\{d_{v}(x)-\bar{d}_{K},0\}$. We also define
$\operatorname{Cols}(\hat{H}_{K},s,n_{H},\epsilon)\in\operatorname{argmax}_{H\subseteq H_{K}\setminus\hat{H}_{K}}\left\{\left|\left|y_{H}\right|\right|_{1}\;\left|\;\;y=((Dx)_{H_{K}}-\bar{d}_{K}\mathbbm{1})_{+},\quad\left|H\right|\leq n_{H},\quad\min_{v\in H}\{y_{v}\}\geq\epsilon\right.\right\},$
which is a subset of the voxels of OAR $K$ with the largest (up to) $n_{H}$ components of $y_{H_{K}\setminus\hat{H}_{K}}$ (those excluded from the current subproblem).
Input : Initial mode $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta)$ with value $z$, number of homogeneity constraints per iteration $n_{S}$, number of OAR constraints per iteration $n_{H}$, thresholds $\tau_{\operatorname{Infs},1}>\tau_{\operatorname{Infs},2}\geq 0,\tau_{z}>0$
Initialize : $m_{S}\leftarrow 1$, $m_{H_{k}}\leftarrow 0,\;\forall k$;
1 while $\sum_{k=1}^{K}m_{H_{k}}+m_{S}>0$ do
2       if $m_{S}>0$ then
3            Solve $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta)$ and obtain its optimal solution $s=(x,\underaccent{\bar}{d},y_{\hat{H}_{K}})$ and optimal value $z$;
4       end if
[…]
6             $z_{\text{prev}}\leftarrow z$;
7             for $k=1,\ldots,K-1$ do
[…]
9             end for
10            Set $\tilde{H}\leftarrow\operatorname{Cols}(s,n_{H},\tau_{\operatorname{Infs},2})$, $m_{H_{K}}\leftarrow|\tilde{H}|$, $\hat{H}_{K}\leftarrow\hat{H}_{K}\cup\tilde{H}$;
11             if $\sum_{k=1}^{K}m_{H_{k}}>0$ then
12                   Solve $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta)$ and obtain its optimal solution $s=(x,\underaccent{\bar}{d},y_{\hat{H}_{K}})$ and optimal value $z$;
13             end if
14
15       end while
[…]
17
18 end while
Output : $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta)$, $s$, $z$
Algorithm 4 Row and column generation algorithm
A.3 Parametric Algorithm
Suppose $s=(\underaccent{\bar}{d},x,y_{\hat{H}_{K}})$ and $s^{\prime}=(\underaccent{\bar}{d}’,x’,y_{\hat{H}_{K}’}’)$ ’ ) are basic feasible solutions, optimal to $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta)$ and $\Phi(S^{\prime},\hat{H}_{1}^{\prime},\ldots,\hat{H}_{K}^{\prime},\beta^{\prime})$, respectively, where without loss of generality, $S\subseteq S^{\prime}$ and $\hat{H}_{k}\subseteq\hat{H}_{k}^{\prime}$, for all $k=1,\ldots,K$. In this context, we say that $s$ and $s’$ ’ are adjacent if once extending subvector $y_{\hat{H}_{K}^{\prime}\setminus\hat{H}_{K}}$ with $y_{\hat{H}_{K}}$ (appending the absent variables), the extended solution $(\underaccent{\bar}{d},x,y_{\hat{H}_{K}’})$ is an adjacent basic feasible solution to $s^{\prime}$ in problem $\Phi(S^{\prime},\hat{H}_{1}^{\prime},\ldots,\hat{H}_{K}^{\prime},\beta^{\prime})$ (that is, each solution can be reached from the other through a single Simplex pivot). We denote this adjacency by the Boolean function $\operatorname{adjacent}(s,s^{\prime})$, returning true if $s$ and $s^{\prime}$ are adjacent, and false otherwise.
Input : Initial bounds $0\leq\beta_{l}<\beta_{u}$, $s^{(1}=s^{(0)}=(\underaccent{\bar}{d}^{(0)},x^{(0)},y_{\hat{H}_{K}}^{(0)})$ optimal for $\beta_{l}$ (and where $\emptyset=\operatorname{Cols}(\hat{H}_{K},s,1,\tau_{\operatorname{Infs},2})$), $\epsilon>0$,
Initialize : $\beta^{(1)}\leftarrow\beta_{l}$, $\beta^{(0)}\leftarrow\beta-\epsilon$.
1 for $\ell=1,\ldots$ do
2      Determine objective coefficient interval $[\beta^{B}_{l},\beta^{B}_{u}]$ that maintains the optimality of current partial basis $B_{s^{(\ell)}}$;
3       if $\left|\left|y_{\hat{H}_{K}}^{(\ell)}\right|\right|_{0}>\theta$ then
4             $\beta^{(\ell)}\leftarrow\min\{\beta^{B}_{u}+\epsilon,\beta_{u}\}$;
5
6       else if $\beta^{(\ell)}>\beta^{(\ell-1)}$ and $s^{(\ell)}$ and $s^{(\ell-1)}$ once extended are adjacent BFS of $\Phi(T,H_{1},\ldots,H_{K})$ then
7            break;
8       else
9            $\beta^{(\ell)}\leftarrow\max\{\beta^{B}_{l}-\epsilon,\beta_{l}\}$ ;
10
11       end if
12      Solve subproblem $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta^{(\ell)})$ to obtain a solution $s^{(\ell+1)}$ and updated constraint sets $(S,\hat{H}_{1},\ldots,\hat{H}_{K})$;
13 end for
Output : $\beta^{(\ell)}$ and solution $s^{(\ell)}$.
Algorithm 5 Parametric penalty algorithm
Appendix B Proofs
B.1 Proof of Lemma 7
Proof.
Since both $(x,\underaccent{\bar}{d},y)$ and $(\tilde{x},\tilde{\underaccent{\bar}{d}},\tilde{y})$ are feasible for problem (13), $(x,\underaccent{\bar}{d},y)\in\mathcal{S}(\beta_{1})$, $\beta_{2}>\beta_{1}$, and $y\geq 0$; thus, the following inequalities hold:
$\underaccent{\bar}{d}-\beta_{1}\sum_{v\in H_{K}}y_{v}\geq\tilde{\underaccent{\bar}{d}}-\beta_{1}\sum_{v\in H_{K}}\tilde{y}_{v}\geq\tilde{\underaccent{\bar}{d}}-\beta_{2}\sum_{v\in H_{K}}\tilde{y}_{v}.$ (14)
Thus, we have obtained the third inequality in the lemma. Additionally, since $(\tilde{x},\tilde{\underaccent{\bar}{d}},\tilde{y})\in\mathcal{S}(\beta_{2})$, the following inequality holds:
$\tilde{\underaccent{\bar}{d}}-\beta_{2}\sum_{v\in H_{K}}\tilde{y}_{v}\geq\underaccent{\bar}{d}-\beta_{2}\sum_{v\in H_{K}}y_{v}.$ (15)
Combining the first inequality of (14) with (15) leads to
$\beta_{1}(\sum_{v\in H_{K}}y_{v}-\sum_{v\in H_{K}}\tilde{y}_{v})\leq\underaccent{\bar}{d}-\tilde{\underaccent{\bar}{d}}\leq\beta_{2}(\sum_{v\in H_{K}}y_{v}-\sum_{v\in H_{K}}\tilde{y}_{v}),$
Which, due to $\beta_{2}>\beta_{1}$, implies that both $\sum_{v\in H_{K}}y_{v}\geq\sum_{v\in H_{K}}\tilde{y}_{v}$ and $\underaccent{\bar}{d}\geq\tilde{\underaccent{\bar}{d}}$. ∎
B.2 Penalty Parameter Bounds Lemma
The next lemma establishes bounds on the value of $\beta^{*}$.
Lemma 9.
Consider an optimal solution of formulation (4) with $\Theta=0$, and let $\lambda_{u}$ be an optimal dual variable value corresponding to constraint (12c). Next, consider an optimal solution of formulation (4) with $\Theta=\lfloor\alpha|H_{K}|\rfloor(\hat{d}_{K}-\bar{d}_{K})+\epsilon$ for every $\epsilon>0$, and let $\lambda_{l}$ be an optimal dual variable value corresponding to constraint (12c). Then, $\lambda_{l}\leq\beta^{*}\leq\lambda_{u}$.
Proof.
Proof.  For the following proofs, we use notation $\mathcal{X}$ to denote the feasible solution set of (13). Also, for every $\Theta\geq 0$, $Z_{C}(\Theta)$ denotes the optimal value of problem (4) with the right-hand side of constraint (12c) given by $\Theta$. Finally, consider the dual function of (4), obtained by dualizing constraint (12c) with dual multiplier $\beta$,
$q(\Theta,\beta)\equiv\beta\Theta+\max_{(x,\underaccent{\bar}{d},y)\in\mathcal{X}}\{\underaccent{\bar}{d}-\beta\sum_{v\in H_{k}}y_{v}\}=\beta\Theta+z(\beta).$
First, to prove the upper bound, let $s^{*}=(x^{*},\underaccent{\bar}{d}^{*},y^{*})$ denote an optimal solution of formulation (4) with $\Theta=0$, and let $\lambda_{u}$ be an optimal dual variable value corresponding to constraint (12c). Strong Lagrangian duality implies that $(x^{*},\underline{d}^{*},y^{*})\in\mathcal{S}(\lambda)$ and $q(0,\lambda)=\lambda\cdot 0+z(\lambda)=\underaccent{\bar}{d}^{*}=Z_{C}(0)$, where the second equality follows from feasibility, which requires $y^{*}=\mathbf{0}$. Evidently, there exists a finite $\lambda$ such that $y(\lambda)=y^{*}=\mathbf{0}$, implying also that $\left|\left|y(\lambda)\right|\right|_{0}=\left|\left|y(\lambda)\right|\right|_{1}=0$. Thus, by Lemma 7, $\beta^{*}\leq\lambda_{u}$.
To prove the lower bound, let $\epsilon>0$, let $s^{*}=(x^{*},\underline{d}^{*},y^{*})$ denote an optimal solution of formulation (4) with $\Theta=\lfloor\alpha|H_{K}|\rfloor(\hat{d}_{K}-\bar{d}_{K})+\epsilon$, and let $\lambda_{l}$ be an optimal dual variable value corresponding to constraint (12c). By definition of $\lambda_{l}$ and from strong duality, we have $(x^{*},\underaccent{\bar}{d}^{*},y^{*})\in\mathcal{S}(\lambda_{l})$, and from the feasibility of $y^{*}$ for problem (4), it follows that $\|y^{*}\|_{1}\leq\Theta$. Moreover, as a result of strong duality,
$q(\Theta,\lambda_{l})=\lambda_{l}\Theta+z(\lambda_{l})=\min_{\beta\geq 0}q(\Theta,\beta)=Z_{C}(\Theta)=\underaccent{\bar}{d}^{*},$
and either $\|y^{*}\|_{1}=\Theta$ or $\lambda_{l}=0$. If $\lambda_{l}=0$, then, by feasibility of $\beta^{*}$ for the dual problem, $\beta^{*}\geq\lambda_{l}=0$. Otherwise (when $\|y^{*}\|_{1}=\Theta$), by definition of $\beta^{*}$, $\left|\left|y(\beta^{*})\right|\right|_{0}\leq\lfloor\alpha\left|H_{K}\right|\rfloor$, which implies that $\left|\left|y(\beta^{*})\right|\right|_{1}\leq\lfloor\alpha\left|H_{K}\right|\rfloor(\hat{d}_{K}-\bar{d}_{K})=\Theta-\epsilon=\|y^{*}\|_{1}-\epsilon$. Thus, it follows from Lemma 7 that $\beta^{*}\geq\lambda_{l}$. ∎
B.3 Proof of Proposition 8
Proof.
Proof.
First observe that at the start of each iteration of Algorithm 5, $\beta>\beta_{\text{prev}}$ unless, in the previous iteration, $\left|\left|y\right|\right|_{0}\leq\lfloor\alpha\left|H_{k}\right|\rfloor$ and $\operatorname{adjacent}(s,s_{\text{prev}})$ does not hold (so a basic feasible solution has been “skipped”). Let $\beta_{i-2},\beta_{i-1},\beta_{i}$ denote a sequence of such penalty parameters $\beta$ in iterations $i-2$, $i-1$ and $i$, respectively, where $\beta_{i-2}<\beta_{i}<\beta_{i-1}$ and where the bases corresponding to $\beta_{i-2}$ and $\beta_{i-1}$ are not adjacent. This implies that once all active constraints have been generated for the basic feasible solution corresponding to $\beta_{i-1}$, a distinct basic feasible solution can be found in $[\beta_{i-2},\beta_{i-1}]$. The finiteness of the number of steps for determining the returned $\beta$ follows from the existence of a finite number of basic feasible solutions. Then, by the uniqueness assumption, the fundamental theorem of LP, and the fact that the algorithm traverses all basic optimal solutions for the penalty parameter values in $[\beta_{l},\beta]$, it follows that the algorithm terminates with $\beta=\beta^{*}$. ∎
Appendix C Conversion of PET Image SUV to OMF
In our experiments, imaging standard uptake values (SUV) are read directly from an FMISO-PET scans. The SUV values are typically normalized with respect to well oxygenated region, and in our case the reference value of $pO2=26$ is fixed as the the median over the entire brain (excluding the tumor) [McKeown, 2014]. These normalized SUV readings are converted to oxygen pressure $pO2$ values using the formula
$pO2(SUV)=\frac{(A-SUV)\times C}{SUV-A+B},$
where $A=10.9$, $B=10.7$ and $C=2.5$ [Toma-Dasu et al., 2012, Lindblom et al., 2017]. Then, the OER is given by
$OER(pO2)=\frac{m\times pO2+K}{pO2+K};$
see for example Titz and Jeraj [2008]. The values assumed here are $K=3$ and $m=3$ following Saka et al. [2014]. Finally, the OMF is simply OER normalized as a proportion of the maximum OER value in the image. Finally, we apply interpolation to deduce voxel OMF values for the CT images that may have a higher resolution of voxels than the corresponding PET images.
Note that the values of the constants in this appendix appear to be highly uncertain according to the scientific literature. In particular, for the conversion of $pO2$ to $OER$Titz and Jeraj [2008] consider a range of possible values $K=2,3,4$. Note that for $pO2=15$ it implies that the corresponding $OMF$ value could vary in the interval $(0.875,0.930)$ (assuming a well oxygenated reference point has $pO2=151$) due only to parameter $K^{\prime}s$ uncertainty, while significant uncertainty may be associated with the values of other parameters as well. Based on such a sensitivity analysis, experimenting with values of $\delta$ in $[0,0.14]$ would be reasonable.
Appendix D Additional Numerical Results
Figure 13 shows the optimal solution value as a function of $\delta$ for different values of $\mu$. In this figure, points that can only be satisfied by the zero beamlet intensity configuration ($x=0$) are omitted from the graph (this applies to all figures throughout the paper). In particular, for each value of $\mu$, only the zero solution is feasible below a certain $\delta$ threshold. For example, $\mu=1.125$ cannot be satisfied with any positive dose (accordingly, positive beamlet intensity) when $\delta\geq 0.03$. This is because $\mu$ constrains the biological homogeneity, which is directly affected by the “size” of the uncertainty set corresponding to the magnitude of the parameters $\delta$ and $\gamma$.
Figure 13: Optimal solution value of (2.2.2) for $\gamma=0.04$ and different values of $\delta$.
Figure 14: Nominal performance of optimal solutions to (2.2.2) with different values of $\gamma$ and $\delta$, compared with the nominal solution.
Figure 15: Physical performance of optimal solutions to (2.2.2) with different values of $\gamma$ and $\delta$, compared with the nominal solution.
Figure 16: Worst-case performance ($\hat{\mu}$, $\hat{\underline{d}}$) of solutions optimal to (2.2.2) with misspecified $\gamma$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\gamma=0.02$ and $\delta$ as indicated. The performance of the nominal solution is also shown for comparison.
Figure 17: Worst-case performance ($\hat{\mu}$, $\hat{\underline{d}}$) of solutions optimal to (2.2.2) with misspecified $\gamma$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\gamma=0.04$ and $\delta$ as indicated. The performance of the nominal solution is also shown for comparison.